### v1

# TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling

Abstract
Large Language Models (LLMs) rely on attention mechanisms whose time complexity grows quadratically with input sequence length, creating significant computational bottlenecks during the prefilling stage. Existing static sparse attention methods typically degrade accuracy, while dynamic sparsity methods introduce additional computational overhead due to runtime sparse index estimation. To address these limitations, we propose TriangleMix, a novel training-free static attention pattern. TriangleMix employs dense attention in shallow layers and switches to a triangle-shaped sparse pattern in deeper layers. Extensive experiments demonstrate that TriangleMix reduces attention overhead by 3.7×3.7$\times$3.7 × to 15.3×15.3$\times$15.3 × in deep layers, and decreases overall Time-to-First-Token (TTFT) by 12% to 32% for sequence lengths ranging from 32K to 128K, without sacrificing model accuracy. Moreover, TriangleMix can be seamlessly integrated with dynamic sparsity methods to achieve further speedup, e.g. accelerating MInference by 19% at 128K, highlighting its potential to enhance LLM inference efficiency. Implementation at https://aka.ms/TriangleMix
TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling
Zhiyuan He1, Yike Zhang2thanks: Work during internship at Microsoft. , Chengruidong Zhang1, Huiqiang Jiang1, Yuqing Yang1, Lili Qiu1 1Microsoft Research, 2Tsinghua University zhiyuhe@microsoft.com
1 Introduction
(a) TriangleMix
(b) TriangleMix combined with dynamic attention
Figure 1: TriangleMix on Llama-3.1-8B-Instruct.
Large Language Models (LLMs) are capable of processing input sequences of varying lengths Grattafiori et al. (2024); Achiam et al. (2023); Yang et al. (2024), a crucial ability that supports diverse downstream tasks, such as question answering Bai et al. (2023), long-form document understanding Zhao et al. (2024), and code generation Jiang et al. (2024b). However, due to the quadratic complexity of attention mechanisms, the time-to-first-token (TTFT) significantly increases as the input context length grows. As prior research has demonstrated, attention computation has become a critical bottleneck in the prefilling stage of LLMs Jiang et al. (2024a); Lai et al. (2025).
To address this bottleneck and accelerate the prefilling stage, researchers have proposed both static and dynamic sparse attention methods. Static sparse attention methods, such as StreamingLLM Xiao et al. (2023), reduce computational complexity from $O(N^{2})$ to $O(N)$ but suffer notable performance degradation on tasks with long contexts Li et al. (2024a). On the other hand, dynamic sparsity approaches, such as MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025), aim to maintain near-full performance by dynamically computing sparse attention indices during inference.
Dynamic sparsity methods are effective for handling extremely long contexts (e.g., beyond 128K tokens). However, they incur additional overhead due to the runtime estimation of sparse block indices. While this overhead is negligible at very long sequence lengths, it becomes non-trivial for moderately long contexts ranging from 32K to 128K tokens. We measured the average attention kernel time per layer on Llama-3.1-8B-Instruct, and the results (Table 1) show that MInference can be 2× slower than dense attention at 32K, and offers only limited speedup at 64K. Similarly, FlexPrefill delivers a modest 1.4× speedup at 32K.
Figure 2: The average gradient $\mathrm{Grad}(\bm{M},l)$ of the Middle Q-K sections, measured on Llama-3.1-8B-Instruct, shows a significant decline in deeper layers. This suggests that the Middle Q-K components in deeper layers contribute minimally and might potentially be skipped to improve efficiency.
In this work, we aim to develop a static attention pattern that effectively accelerates the prefilling stage across various context lengths. To achieve this goal, we propose a novel gradient-based method to identify the crucial sections within attention maps. Our analysis demonstrates that the middle Q-K section (Figure 2 and 3) of the attention map can be safely omitted in deeper layers without incurring performance loss. Motivated by this insight, we introduce TriangleMix, a lossless and efficient static attention pattern for accelerating long-context LLMs. As depicted in Figure 1, TriangleMix applies standard dense attention in shallow layers, and switches to a triangle-shaped sparse attention pattern in deeper layers. TriangleMix offers four major advantages:
  • It is training-free, enabling direct integration with state-of-the-art pretrained LLMs.
  • It achieves nearly lossless accuracy, preserving the full capabilities of the underlying models.
  • Its static triangle-shaped sparsity eliminates the need for dynamic block index estimation and reduces the computational complexity from $O(N^{2})$ to $O(N)$.
  • It is complementary to dynamic attention approaches. Replacing dynamic attention in deep layers with the static Triangle attention further enhances acceleration, compared with dynamic sparse attention alone.
We conduct extensive experiments on three long-context LLMs, including Llama-3.1-8B-Instruct Grattafiori et al. (2024), Llama-3-8B-Instruct-262K GradientAI (2024), and Qwen2.5-7B-Instruct Yang et al. (2024), using all tasks from the RULER and LongBench benchmarks Hsieh et al. (2024); Bai et al. (2023). The results demonstrate that TriangleMix maintains the same accuracy as full attention. For Llama-3.1-8B-Instruct, TriangleMix decreases overall TTFT by 12% to 32% across context lengths ranging from 32K to 128K. Moreover, it can be seamlessly combined with dynamic sparsity methods without degrading performance. For example, integrating TriangleMix with MInference decreases TTFT by 19% at 128K context length for Llama-3.1-8B-Instruct.
2 Methodology
2.1 Sparse Attention
Figure 3: Three attention sections.
Figure 4: First row: Average attention score $\mathrm{Att}(\bm{M},l)$ for the Middle and Last Q-K sections; Second row: average gradient $\mathrm{Grad}(\bm{M},l)$ for the Middle and Last Q-K sections.
Method 32K 64K 128K
Dense 44 179 750
MInference 94 152 229
FlexPrefill ($\gamma=0.95$) 32 71 220
Table 1: Averaged attention kernel time in milliseconds for one layer on Llama-3.1-8B-Instruct.
The prefilling stage of Transformer attention can be formulated as:
$\bm{A}=\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M}))$
[…]
To accelerate computing, sparse attention aims to find a sparse mask matrix $\bm{M}^{\prime}$ to compute the attention output:
$\bm{A^{\prime}}=\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M^{\prime}}))$
where $|\bm{A}-\bm{A}^{\prime}|$ is expected to be sufficiently small to preserve the original model performance.
Existing sparse attention methods fall into two categories: static sparsity and dynamic sparsity. In static sparsity, the sparse mask $\bm{M}^{\prime}$ is fixed and does not depend on the input. For example, StreamingLLM shows that most attention scores accumulate on the first few sink tokens and nearby tokens in a sliding window Xiao et al. (2023). This leads to the following streaming pattern:
$\bm{M}^{\mathrm{streaming}}_{i,j}=\begin{cases}1,&i\geq j,j\leq si\\ 1,&i\geq j,i-j\leq sl\\ 0,&\mathrm{otherwise}\end{cases}$
where $si$ is the number of sink tokens and $sl$ is the sliding window size. This pattern, also called "A-shape" in previous works Jiang et al. (2024a), is very efficient with only $O(N)$ complexity. However, it causes significant performance drops for long-context LLMs Li et al. (2024a). In contrast, dynamic sparsity methods, such as MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025), estimate $\bm{M}^{\prime}$ at inference time based on the input through an observation window or the pooling vector of $\bm{Q}$ and $\bm{K}$.
Dynamic sparsity generally provides better accuracy and acceleration for extremely long contexts (e.g., contexts exceeding 128K tokens). However, for moderately long contexts, the online estimation introduces additional time overhead. Specifically, as demonstrated in Table 1, MInference is slower than dense attention or offers small acceleration for contexts below 64K, while FlexPrefill achieves only modest improvements (approximately 1.4x speedup for 32K context).
2.2 Attention Sections
In this paper, we aim to design a static sparse attention mask that is both efficient and effective. Although a large fraction of attention scores falls into the streaming pattern, a significant performance gap still remains between streaming sparse attention and full dense attention. This leads us to the following question:
Beyond the streaming pattern, what other parts of the attention are crucial for maintaining the end-to-end performance of long-context LLMs?
To answer this, we divide the full attention matrix into three distinct sections as illustrated in Figure 3:
  • Streaming section: includes attention sink and the sliding window;
  • Last Q-K section: covers interactions between the last part of $\bm{Q}$ and $\bm{K}$, excluding the Streaming section;
  • Middle Q-K section: consists of the remaining interactions between the middle parts of $\bm{Q}$ and $\bm{K}$.
We define the Last Q-K mask as:
$\bm{M}^{\mathrm{last}}_{i,j}=\begin{cases}1,&i\geq j,N-i<last,\\ &j>si,i-j>sl\\ 0,&\mathrm{otherwise}\end{cases}$
[…]
Similarly, the Middle Q-K mask is defined as:
$\bm{M}^{\mathrm{middle}}_{i,j}=\begin{cases}1,&i\geq j,N-i\geq last,\\ &j>si,i-j>sl\\ 0,&\mathrm{otherwise}\end{cases}$
2.3 Probing Attention Importance
Firstly, we inspect the average attention scores within the two sections defined by masks $\bm{M}^{\mathrm{middle}}$ and $\bm{M}^{\mathrm{last}}$. We observe that these average attention scores are quite similar, although both are notably lower than the average score observed in the streaming section $\bm{M}^{\mathrm{streaming}}$. To achieve a more precise measurement of the relative importance of these sections, we introduce a novel gradient-based probing technique.
Specifically, given an input $\bm{X}_{\mathrm{input}}$ containing $N$ tokens fed into a language model, we define a probing variable $\bm{\theta}\in\mathbb{R}^{L\times N\times N}$, where $\bm{\theta}_{\ell,i,j}=1$ for all layers $\ell$ and token indices $i,j$. Here, $L$ represents the total number of layers in the model. For simplicity of notation, we omit the attention head dimension here, but our proposed method generalizes naturally to both Multi-head Attention Vaswani et al. (2017) and Grouped Query Attention Ainslie et al. (2023).
In layer $l$, the attention scores is calculated as:
$\bm{A_{\theta}}=\bm{\theta}_{l}\odot\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M})\right)$
where $\odot$ denotes element-wise multiplication. Since all elements of $\bm{\theta}$ are initially set to 1, this operation does not alter the attention scores.
The model then outputs a logit prediction $\bm{Y_{\theta}}$ for the next token, represented as a vector $\bm{Y_{\theta}}=[y_{1},y_{2},\dots,y_{\text{vocab}}]$. Our focus is on a specific logit $y_{\mathrm{gt}}$, associated with the correct label or ground truth token. For instance, the correct answer in a multiple-choice scenario or the initial token of the target sequence in a needle-in-a-haystack task.
We are particularly interested in the partial derivative $\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}$, which measures the sensitivity of the model’s output to changes in attention scores. We quantify the importance of a specific attention section by computing the mean of the derivative values within that section. Formally, we define the gradient-based importance as:
$\mathrm{Grad}(\bm{M},l)=\frac{\sum_{\bm{M}_{i,j}=1}{\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}}}{\sum_{\bm{M}_{i,j}=1}{\bm{M}_{i,j}}}$
where $\bm{M}$ denotes the binary mask representing the targeted attention section, and $\mathrm{Grad}(\bm{M},l)$ indicates the average sensitivity of $y_{\mathrm{gt}}$ to changes within this section at layer $l$.
For comparison, we also define the average attention score of the section as:
$\mathrm{Att}(\bm{M},l)=\frac{\sum_{\bm{M}_{i,j}=1}{\bm{A}_{l,i,j}}}{\sum_{\bm{M}_{i,j}=1}{\bm{M}_{i,j}}}$
[…]
Empirically, we leverage a key-value retrieval task to analyze the importance of different sections of attention. The language model is tasked with retrieve the corresponding value, starting from its first generated token. $y_{gt}$ is set to be the first token of the correct value. The full task prompt is provided in Appendix A.1.
We calculate $\mathrm{Grad}(\bm{M},l)$ and $\mathrm{Att}(\bm{M},l)$ for both the middle Q-K section ($\bm{M}^{\mathrm{middle}}$) and the last Q-K section ($\bm{M}^{\mathrm{last}}$) across 100 randomly generated samples. The hyperparameters are set as follows: the number of sink tokens ($si$) is 64, the sliding window size ($sl$) is 128, and the length of the last attention section ($last$) is also set to 128. Each input sequence consists of approximately 2,000 tokens. Finally, we visualize the results for $\mathrm{Grad}(\bm{M},l)$ and $\mathrm{Att}(\bm{M},l)$ in Figure 4. For clarity, attention scores and gradient values are both summed across all heads within each layer.
2.4 TriangleMix
Methods Average
2Wiki
GovRep
Hotpot
LCC
MNews
MF-en
PsgCnt
PsgRtr
Qasper
Rbench
Samsum
TREC
TrvQA
Llama-3.1-8B-Instruct 54.8 48.1 34.4 61.6 66.6 25.6 56.9 16.8 99.7 44.9 52.7 42.6 71.0 91.6
Streaming 34.7 17.4 32.2 19.8 65.6 24.6 25.6 5.7 11.8 23.3 54.1 40.6 52.7 77.3
Triangle 47.0 40.6 33.2 52.3 63.6 24.6 53.5 7.9 42.0 46.3 46.5 43.0 67.0 90.1
StreamingMix 48.9 36.4 34.3 39.4 64.9 25.8 42.3 18.4 98.3 42.4 53.6 42.6 51.0 86.8
DuoAttention 54.5 44.6 34.1 59.1 68.0 25.3 55.0 15.7 99.7 43.9 58.8 42.7 71.3 91.0
[…]
Llama-3-8B-Instruct-262k 44.2 21.0 34.3 26.4 46.1 26.3 50.2 4.7 95.0 30.5 43.3 41.1 68.3 86.8
Streaming 30.8 16.8 34.3 18.5 41.8 25.3 39.6 0.3 10.0 23.4 36.5 38.6 49.0 66.0
Triangle 36.2 17.1 34.1 21.2 35.5 25.3 50.0 5.0 32.0 28.9 32.4 40.5 63.0 85.9
StreamingMix 39.9 18.5 34.4 19.8 42.3 25.9 47.4 5.3 85.0 30.0 38.3 40.1 48.3 83.3
DuoAttention 42.6 19.2 34.6 25.1 44.4 26.1 51.8 2.3 87.7 29.6 41.1 40.4 65.7 85.9
[…]
Qwen2.5-7B-Instruct 46.8 24.5 30.7 32.2 62.2 22.3 41.3 13.1 99.3 24.4 60.7 42.5 67.0 88.7
Streaming 28.3 10.1 31.3 9.5 61.7 21.5 20.7 6.5 8.2 11.4 44.8 40.2 51.7 49.9
Triangle 36.9 17.4 27.5 24.0 48.5 17.9 38.5 5.6 43.9 22.3 39.6 41.2 65.0 88.1
StreamingMix 45.2 24.4 31.9 38.5 68.2 22.5 43.5 10.7 93.0 24.8 60.9 42.0 43.7 83.1
[…]
Table 2: Comparison with static sparsity methods on Longbench. TriangleMix is our proposed method. StreamingMix is a hybrid baseline that uses dense attention in shallow layers but replaces Triangle attention with a streaming pattern in deeper layers. The number of dense layers matches that of TriangleMix.
From Figure 4, we observe that both the Middle Q-K and Last Q-K sections have low average attention scores, with the Middle Q-K section slightly higher across all layers. However, when examining the gradient values, we find that the derivatives in the Last Q-K section are significantly larger than those in the Middle Q-K section. This indicates that the model output is more sensitive to changes in the attention scores of the Last Q-K section compared to the Middle Q-K section.
Besides, for the last layer, we observe that we always have $\mathrm{Grad}(\bm{M},l)=0$. This phenomenon is expected due to the specific structure of the attention mechanism. In the last layer, the logit corresponding to the first predicted token is computed solely based on the last query vector and the entire key matrix $\bm{K}^{T}$. Formally, this corresponds to computing only the last row of the attention map, represented as $\mathrm{Softmax}(\bm{Q}_{\mathrm{last}}\bm{K}^{T})$. Since the gradient measure $\mathrm{Grad}(\bm{M},l)$ is defined based on the derivative of the ground truth output logit with respect to parameters associated with the attention sections defined by the mask, and because the first output token logit does not depend on the Middle Q-K sections in the last layer (which involve query vectors from other positions), the resulting gradient naturally becomes zero.
During the training phase, this observation does not apply. The training loss is typically computed using all output logits across the entire input sequence, which requires calculating the full attention matrix in the last layer. However, during inference, since our primary interest lies only in the prediction, we can optimize the computation in the final layer by restricting attention calculation to $\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{Q}_{\mathrm{last}}\bm{K}^{T})$.
Moreover, we observe that the gradient values for the Middle Q-K sections follow a consistent trend across all models: they remain moderate in the shallow layers but drop sharply in deeper layers. As shown in Figure 2, this pattern is evident. For Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, the decline occurs between layers 10 and 20, while for Qwen2.5-7B-Instruct, it starts after layer 20.
This suggests that the attention from the Middle Q-K sections in deep layers contribute less to the output logits. Notably, the number of elements in the Middle Q-K sections scales quadratically with the input token length $N$, whereas the elements in the Streaming and Last Q-K sections grow only linearly with $N$. Therefore, if the computation of the Middle Q-K sections is skipped in deep layers, the time complexity of attention can be reduced from $O(N^{2})$ to $O(N)$. And we assume such skipping will not cause significant loss in output quality.
Based on these observations, we propose TriangleMix, an effective and efficient static attention pattern designed for long-context LLMs. We denote the starting layer of Triangle attention as $L_{\mathrm{tri\_start}}$. For layer $l\leq L_{\mathrm{tri\_start}}$, we apply standard dense attention:
$\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M})\right)$
where $\bm{M}$ is the standard causal attention mask.
In layer $l>L_{\mathrm{tri\_start}}$, we skip the Middle Q-K sections and apply a Triangle-shape attention pattern:
$\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c\left(1-(\bm{M}-\bm{M}^{\text{middle}})\right)\right)$
where $\bm{M}^{\text{middle}}$ represents the mask corresponding to the Middle Q-K sections.
Thus, in deep layers, the attention complexity becomes $O(N)$, and there is no need for dynamic estimation of the attention mask, making the computation highly efficient. Furthermore, we can optionally adopt dynamic sparse attention in shallow layers to further accelerate computation, while keeping the static Triangle pattern in deep layers.
We acknowledge that the Triangle-shape attention pattern was first introduced in Scbench Li et al. (2024a). However, Scbench applies the Triangle-shape attention uniformly across all layers, leading to notable performance degradation. In contrast, by using a gradient-based probing method, we find that the Middle Q-K sections remain important in shallow layers. Therefore, by combining dense (or dynamic sparse) attention in shallow layers and Triangle-shape attention in deep layers, TriangleMix preserves model performance while achieving significant efficiency gains.
3 Evaluations
3.1 Settings
LLMs. We evaluate our method on the following long-context language models. (1) Llama-3.1-8B-Instruct, a state-of-the-art model from Meta with a context window of up to 128K tokens Grattafiori et al. (2024). (2) Llama-3-8B-Instruct-262K, a fine-tuned variant from Llama-3-8B-Instruct that supports up to 262K tokens GradientAI (2024). (3) Qwen2.5-7B-Instruct, a model from Alibaba with a maximum context window of 128K Yang et al. (2024).
Static Sparsity Baselines. We compare our method with four static sparsity baselines. The first is the Streaming pattern Xiao et al. (2023), applied to all layers with $si=8$ and $sl=512$. The second is Triangle Attention applied to all layers, using $si=8$, $sl=512$, and $last=128$. The third is a hybrid baseline StremingMix, where dense attention is used for layers $l\leq L_{\text{tri\_start}}$, and streaming attention is used for deeper layers.
We also compare against DuoAttention, which learns a separate sparsity pattern for each attention head Xiao et al. (2024). DuoAttention provides a trained pattern for Llama-3.1-8B-Instruct; we train a sparsity pattern for Llama-3-8B-Instruct-262K using the authors’ script. However, for Qwen2.5-7B-Instruct, the training does not converge, so its result is omitted. At the prefilling stage, half of the heads follow the streaming pattern, while the other half use dense attention.
Dynamic Sparsity Baselines. We compare with two dynamic sparsity methods: MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025). We apply default sparsity settings of all LLMs from MInference. For FlexPrefill, we test its hyperparameter $\gamma=0.90$ and $\gamma=0.95$. Additionally, our method can be combined with dynamic sparsity attention: Dynamic sparsity (MInference or FlexPrefill) is applied for layers $l\leq L_{\text{tri\_start}}$, while Triangle Attention is used for deeper layers. This combination reduces runtime overhead because Triangle Attention avoids runtime mask estimation in dynamic sparse attention. We use “Ours + MInference” and “Ours + FlexPrefill” to denote the corresponding combined versions.
Hyperparameters. The hyperparameter $L_{\text{tri\_start}}$ determines the layer from which the Triangle Attention pattern is applied. For layers $l\leq L_{\text{tri\_start}}$, we use dense attention or dynamic sparsity methods. For layers $l>L_{\text{tri\_start}}$, we switch to the static Triangle Attention. Based on the observations in Figure 4, we set $L_{\text{tri\_start}}=16$ for Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, and $L_{\text{tri\_start}}=20$ for Qwen2.5-7B-Instruct. For the Triangle mask, we use: sink tokens $si=8$, sliding window size $sl=512$, and last window size $last=128$, which is the same with the sparse sparsity baselines.
Benchmarks. We evaluate on two benchmarks: (1) RULER: A synthetic long-context benchmark including tasks like hay-in-the-stack, KV retrieval, variable tracking, and QA. We evaluate at input lengths of 4K, 8K, 16K, 32K, 64K, and 128K, with 100 samples per task-length pair Hsieh et al. (2024). (2) LongBench: A realistic long-context benchmark with tasks such as single-doc and multi-doc QA, few-shot learning, code generation, retrieval, summarization, and counting Bai et al. (2023). We evaluate all English tasks.
Specifically, Qwen2.5-7B-Instruct adopts Yarn scaling to improve its long context ability for input length larger than 32K Peng et al. (2023). We report the performance of the origin Qwen2.5-7B-Instruct model for RULER length < 32K, and the Yarn-scaled model for RULER length 32K, 64K and 128K. For LongBench, we only evaluate the origin Qwen2.5-7B-Instruct model for simplicity.
3.2 Effectiveness of TriangleMix
Methods 4K 8K 16K 32K 64K 128K Avg.
Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
Streaming 64.1 55.4 40.5 28.9 26.7 3.3 36.5
Triangle 88.8 88.3 82.8 72.6 65.0 39.0 72.7
StreamingMix 94.3 91.8 90.2 86.2 79.2 71.9 85.6
DuoAttention 95.7 93.1 88.4 84.3 82.5 64.0 84.7
[…]
Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
Streaming 49.4 38.7 33.4 30.2 26.5 21.7 33.3
Triangle 88.1 84.7 80.6 71.0 65.0 55.4 74.1
StreamingMix 87.3 83.2 80.8 79.9 77.6 70.8 79.9
DuoAttention 92.8 91.9 89.0 85.0 81.1 76.1 86.0
[…]
Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
Streaming 55.6 47.8 35.9 30.6 28.1 21.0 36.5
Triangle 88.4 81.9 74.0 58.3 51.0 14.6 61.4
StreamingMix 93.3 90.4 89.4 81.7 76.4 63.6 82.5
[…]
Table 3: Comparison with static sparsity methods on RULER. Llama-3.1, Llama-3-262K, Qwen2.5 are abbreviations for Llama-3.1-8B-Instruct, Llama-3-8B-262K, and Qwen2.5-8B-Instruct. The same applies to Table 4.
Comparison with Static Sparsity. Table 3 and Table 2 present the evaluation results of TriangleMix and various static sparsity baselines. TriangleMix consistently outperforms all static baselines and remains nearly lossless performance across all input lengths. In contrast, other static methods usually suffer noticeable performance degradation. DuoAttention performs comparably to TriangleMix on LongBench using Llama-3.1-8B-Instruct. However, its performance is worse than TriangleMix on RULER, especially deteriorates at longer lengths (e.g., 64K and 128K). Furthermore, DuoAttention requires a separate training phase to learn head-wise sparsity patterns, which introduces additional computational overhead and may fail to converge on models such as Qwen-2.5-7B-Instruct.
On the LongBench benchmark, we observe that the full Triangle pattern performs poorly on PassageRetrieval tasks, while the StreamingMix pattern underperforms on the in-context learning task (TREC). These findings suggest that attention over the Middle Q-K section in shallow layers is crucial for retrieval tasks, while attention over the Last Q-K section in deeper layers plays a key role in supporting in-context learning.
Comparison with Dynamic Sparsity. Table 4 and Table 8 report results comparing TriangleMix with dynamic sparsity methods, as well as their combinations. FlexPrefill when $\gamma=0.90$ usually leads to performance degradation, and MInference generally outperforms it. While the performance of MInference and FlexPrefill with $\gamma=0.95$ is similar on the RULER benchmark, MInference still achieves better results on LongBench. When combined with either MInference or FlexPrefill, TriangleMix always fully retains the original performance of the dynamic sparsity methods. Moreover, combining TriangleMix with FlexPrefill even yields slight performance gains, particularly with Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct.
Overall, TriangleMix preserves the full accuracy of dense attention across tasks while significantly reducing time complexity from $O(N^{2})$ to $O(N)$ for layers $l>L_{\text{tri\_start}}$.
Methods 4K 8K 16K 32K 64K 128K Avg.
Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
MInference 96.3 95.1 95.0 90.5 86.8 75.0 89.8
[…]
FlexPrefill (0.90) 90.1 89.7 89.4 89.1 82.0 74.1 85.7
[…]
FlexPrefill (0.95) 95.0 94.7 94.5 92.8 86.5 75.9 89.9
[…]
Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
MInference 93.4 90.7 88.8 85.1 82.6 80.1 86.8
[…]
FlexPrefill (0.90) 82.4 80.0 80.7 76.2 74.5 69.1 77.1
[…]
FlexPrefill (0.95) 90.3 87.7 88.0 83.8 80.2 75.7 84.3
[…]
Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
MInference 95.6 93.8 92.8 84.9 79.1 68.0 85.7
[…]
FlexPrefill (0.90) 84.4 79.4 82.4 71.4 65.4 47.7 71.8
[…]
FlexPrefill (0.95) 91.1 89.7 88.5 75.4 72.5 51.2 78.1
[…]
Table 4: Comparison with dynamic sparsity methods on RULER.
3.3 Analysis of Start Layer
Figure 5: Average RULER score at 64K length for different $L_{\mathrm{tri\_start}}$ values.
In this section, we investigate the choice of $L_{\text{tri\_start}}$ for each LLM. We evaluate all models on the 64K length tasks in the RULER benchmark, sweeping $L_{\text{tri\_start}}$ from 0 to the total number of layers with a step size of 2. As shown in Figure 5, Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit similar patterns: setting $L_{\text{tri\_start}}=12$ retains 99.7% and 100.0% of their original performance, respectively. This indicates that 62.5% of the layers can adopt the $O(N)$ Triangle Attention without noticeable degradation. In contrast, Qwen-2.5-7B-Instruct requires a higher threshold: setting $L_{\text{tri\_start}}=20$ retains 97.9% of the performance, corresponding to a sparsity ratio of 28.6%.
Method 32K 64K 128K
Dense 44 179 750
MInference 94 (0.5x) 152 (1.2x) 229 (3.3x)
FlexPrefill (0.95) 32 (1.4x) 71  (2.5x) 220 (3.4x)
FlexPrefill (0.90) 24 (1.8x) 52  (3.4x) 143 (5.2x)
[…]
Table 5: Attention kernel latency (ms) for one layer measured on Llama-3.1-8B-Instruct.
Figure 6: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct.
Method 32K 48K 64K 80K 96K 112K 128K
Dense 4.1 7.3 11.2 15.9 21.3 27.5 34.5
MInference 5.5 (+34%) 7.8 (+7%) 10.1 (-10%) 12.3 (-23%) 13.4 (-37%) 15.9 (-42%) 18.0 (-48%)
FlexPrefill (0.95) 3.6 (-12%) 5.5 (-25%) 7.7 (-31%) 9.9 (-38%) 12.3 (-42%) 15.0 (-45%) 17.8 (-48%)
TriangleMix 3.6 (-12%) 5.9 (-19%) 8.6 (-23%) 11.7 (-26%) 15.2 (-29%) 19.1 (-31%) 23.4 (-32%)
Ours + MInference 4.2 (+2%) 6.0 (-18%) 7.7 (-31%) 9.5 (-40%) 10.9 (-49%) 12.7 (-54%) 14.5 (-58%)
Ours + FP (0.95) 3.4 (-17%) 5.2 (-29%) 7.2 (-36%) 9.2 (-42%) 11.3 (-47%) 13.6 (-51%) 16.1 (-53%)
FlexPrefill (0.90) 3.3 (-20%) 5.1 (-30%) 6.8 (-39%) 8.9 (-44%) 11.1 (-48%) 13.6 (-51%) 15.5 (-55%)
Ours + FP (0.90) 3.3 (-20%) 4.5 (-38%) 6.7 (-40%) 8.5 (-47%) 10.5 (-51%) 12.6 (-54%) 14.5 (-58%)
Table 6: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct. We separate FlexPrefill ($\gamma=0.90$) as it generally leads to accuracy degradation, whereas the other settings do not.
3.4 Efficiency of TriangleMix
We implement Triangle Attention using Triton Tillet et al. (2019), with further implementation details provided in Appendix A.2. All experiments are conducted on a single NVIDIA A100 80GB GPU. Dense attention is implemented using FlashAttention Kwon et al. (2023).
Attention Kernel Latency. We first benchmark the average attention kernel time per layer on Llama-3.1-8B-Instruct at sequence lengths of 32K, 64K, and 128K. As shown in Table 5, Triangle Attention significantly outperforms dense attention, achieving 3.7×3.7$\times$3.7 × to 15.3×15.3$\times$15.3 × speedup. The speedup mainly comes from Triangle Attention’s linear complexity ($O(N)$), which makes it much more efficient than $O(N^{2})$ dense attention. Besides, unlike dynamic sparsity methods, it doesn’t require estimating attention block indices during runtime, which saves additional overhead.
End-to-End TTFT. We further measure time-to-first-token (TTFT) across sequence lengths of 32K, 48K, 64K, 96K, 112K, and 128K. Results for Llama-3.1-8B-Instruct are reported in Table 6. TriangleMix applies dense attention in shallow layers and Triangle Attention in deeper layers, reducing TTFT by 12%–32%. It is also compatible with dynamic attention for enhanced performance. For example, integrating MInference with TriangleMix reduces TTFT from 18.0s to 14.5s (a 19% decrease). Under the lossless setting, TriangleMix combined with FlexPrefill ($\gamma=0.95$) achieves the fastest TTFT for 32K–80K sequences, while TriangleMix with MInference is optimal for 96K–128K. We visualize the results of some selected methods in Figure 6.
4 Discussion
In this section, we discuss why Middle Q-K attention can be safely removed. Our hypothesis is that the observed phenomena arise from a Train–Test Misalignment. During training, the loss is applied uniformly to all positions in the input sequence. For example, if the input contains 2048 tokens, the model is also trained to predict token 1001 given tokens 1–1000. However, during inference, the primary goal is to predict tokens after the prompt (such as token 2049), making predictions of tokens inside the prompt (such as token 1001) largely irrelevant.
To validate this assumption, we conduct an experiment on Llama-3.1-8B-Instruct. We sample 100 Wikipedia articles and extract the first 2048 tokens of each. In the last 16 layers, we apply Triangle Attention, skipping attention for Q indices between 512 and 1024. We then measure changes of the average perplexity (PPL) relative to dense attention in two output spans.
Method PPL (1024–1152) PPL (1920–2048)
Dense 8.02 7.79
TriangleMix 8.34 (+0.32) 7.81 (+0.02)
Table 7: With 2048 input tokens, we measure perplexity on two output spans. Triangle attention is applied, skipping attention between Q indices 512–1024 in the last 16 layers. Experiments are on Llama-3.1-8B-Instruct.
The results in Table 7 show that skipping attention of Q index 512–1024 in deep layers noticeably increases perplexity when predicting tokens in the 1024–1152 range. For these positions, the skipped attention effectively becomes the Last Q-K attention necessary for predicting the next token. In contrast, for positions beyond 1920, the effect is minimal.
The Middle Q-K region is important primarily for predicting tokens within the prompt, while for generation tasks focused on tokens after the prompt, the Middle Q-K in deeper layers is largely redundant. Although further theoretical work is needed, this observation suggests that selective sparsification of attention during inference can be done safely.
5 Related Works
Static Sparsity Attention. Early methods employ fixed sparse patterns, such as strided Child et al. (2019), dilated Ding et al. (2023), sliding window Jiang et al. (2023), and mixed patterns Beltagy et al. (2020), typically requiring training from scratch. DuoAttention Xiao et al. (2024) introduces head-level static sparsity but needs additional offline training. Training-free approaches like Streaming Xiao et al. (2023) and Triangle Li et al. (2024a) attention often degrade accuracy for long contexts. In contrast, the proposed TriangleMix attention pattern significantly mitigates performance loss and is nearly lossless in accuracy.
Dynamic Sparsity Attention. Methods like MInference Jiang et al. (2024a), FlexPrefill Lai et al. (2025), and XAttention Xu et al. (2025) dynamically determine attention masks online, maintaining high accuracy but incurring computational overhead for moderately long contexts. TriangleMix can enhance these dynamic approaches for further acceleration .
Long-context LLM Inference. FlashAttention Dao et al. (2022) speeds up attention by reducing memory access through fused operations. PagedAttention Kwon et al. (2023) improves decoding by managing KV cache allocation efficiently. KV cache optimizations techniques also include token-level eviction (SnapKV Li et al. (2024b)) and query-aware cache selection (Quest Tang et al. (2024)). TriangleMix is orthogonal to these approaches.
6 Conclusion
In this paper, we propose a hybrid attention pattern for long-context LLMs, where dense attention is used in shallow layers and Triangle attention is applied in deeper layers. This pattern is training-free and achieves nearly lossless accuracy. Our method achieves a 12% to 32% time-to-first-token reduction across 32K to 128K inputs. Moreover, it can be further accelerated when combined with dynamic sparsity attention.
Limitations
One limitation of our method is that the sparsity ratio is fixed within the pretrained model. We observe that Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit higher inherent sparsity compared to Qwen2.5-7B-Instruct, which constrains the achievable acceleration on the latter. It may be possible to enhance the sparsity of models like Qwen2.5-7B-Instruct through a post-training optimization process, which we leave for future exploration.
Appendix A Appendix
A.1 Task to Probe Attention Importance
We use a key-value retrieval task to analyze the importance of different attention sections. The language model is prompted to retrieve the correct value, beginning from its first generated token. The prompt is as follows:
[…]
We randomly generate pairs of keys and values using Version 4 UUID strings.
A.2 Implementation of Triangle Attention
We implement Triangle Attention using Triton Tillet et al. (2019), with details provided in Algorithm 1. For each row index, we apply a different attention mechanism: if the row index is less than $N-N_{\text{last}}$, we apply Streaming Attention; otherwise, we apply chunked FlashAttention to increase GPU utilization. The outputs from both segments are then merged to produce the final result.
Algorithm 1 Fused Triangle Attention
Input data: $\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V}\in\mathbb{R}^{N\times d_{h}}$
Input triangle shape: sink token number $N_{\mathrm{sink}}$, sliding window size $N_{\mathrm{window}}$, last rows number $N_{\mathrm{last}}$
Input kernel block shape: $B_{M},B_{N}$
 Calculate best number of splits $S$ for last row
 Initialize $\boldsymbol{O}\leftarrow(0)^{N\times d_{h}}$
 Initialize output buffer $\boldsymbol{O}_{\mathrm{last}}\leftarrow(0)^{SN_{\mathrm{last}}\times d_{h}}$
 Initialize LSE buffer $\boldsymbol{l}_{\mathrm{last}}\leftarrow(-\inf)^{SN_{\mathrm{last}}}$
Parallelized in GPU
for $i\leftarrow 1$ to $\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil+S\lceil\frac{N_{\mathrm{last}}}{B_{M}}\rceil$ do
  Initialize $\boldsymbol{O}_{\text{chip}}\leftarrow(0)^{B_{M}\times d_{h}}$
  Initialize $\boldsymbol{m}\leftarrow(-\inf)^{B_{N}}$
  Initialize $\boldsymbol{s}\leftarrow(0)^{B_{M}}$
  if $i\leq\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil$ then
    Upper part: the same as streaming attention
   Load $\boldsymbol{Q}_{\text{chip}}\leftarrow\boldsymbol{Q}^{iB_{M}:(i+1)B_{M}}$
    Loop through sink tokens
   for $j\leftarrow 1$ to $\lceil\frac{N_{\mathrm{sink}}}{B_{N}}\rceil$ do
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
    $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
   end for
    Loop through sliding window
   for $j\leftarrow\lceil\frac{N-N_{\mathrm{window}}}{B_{N}}\rceil$ to $\lceil\frac{N}{B_{N}}\rceil$ do
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
    $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
   end for
    Write outputs
   Save $\boldsymbol{O}^{iB_{M}:(i+1)B_{M}}\leftarrow\boldsymbol{O}_{\text{chip}}$
  else
    Last rows: split in to $S$ chunks
   Chunk index $c\leftarrow\lfloor(i-\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil)/{\lceil\frac{N_{\mathrm{last}}}{B_{M}}\rceil}\rfloor$
   Chunk offset $b\leftarrow(i-\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil)\mod{\lceil\frac{N_{\mathrm{last}}}{B_{M}}\rceil}$
   Q index $i_{Q}\leftarrow\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil+b$
   Load $\boldsymbol{Q}_{\text{chip}}\leftarrow\boldsymbol{Q}^{i_{Q}B_{M}:(i_{Q}+1)B_{M}}$
   for $j\leftarrow c\lceil\frac{N}{SB_{N}}\rceil$ to $(c+1)\lceil\frac{N}{SB_{N}}\rceil$ do
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
    $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
   end for
    Write outputs
   Save $\boldsymbol{O}_{\mathrm{last}}^{(i-i_{Q})B_{M}:(i-i_{Q}+1)B_{M}}\leftarrow\boldsymbol{O}_{\text{chip}}$
   Save $\boldsymbol{l}_{\mathrm{last}}^{(i-i_{Q})B_{M}:(i-i_{Q}+1)B_{M}}\leftarrow\ln\boldsymbol{s}+\boldsymbol{m}$
  end
end for
Merge last row output buffer
$\boldsymbol{O}^{(N-N_{\mathrm{last}}):N}\leftarrow\mathrm{merge\_output}(\boldsymbol{O}_{\mathrm{last}},\boldsymbol{l}_{\mathrm{last}})$
A.3 Additional Longbench Data
Table 8 compares our method with dynamic sparsity attention on Longbench. When combined with MInference, our method achieves similar performance to MInference alone. Integrating with FlexPrefill yields a slight improvement over its original results.
Methods Average
2Wiki
GovRep
Hotpot
LCC
MNews
MF-en
PsgCnt
PsgRtr
Qasper
Rbench
Samsum
TREC
TrvQA
Llama-3.1-8B-Instruct 54.8 48.1 34.4 61.6 66.6 25.6 56.9 16.8 99.7 44.9 52.7 42.6 71.0 91.6
MInference 54.9 48.1 34.5 61.4 67.0 25.9 57.0 17.5 99.7 44.7 52.5 42.6 71.3 91.9
[…]
FlexPrefill; $\gamma=0.90$ 44.0 31.0 32.8 48.8 59.1 25.1 51.3 4.0 39.0 37.1 50.8 41.5 65.3 86.2
[…]
FlexPrefill; $\gamma=0.95$ 48.4 39.4 33.7 58.3 67.2 25.6 55.6 4.0 47.7 42.5 53.8 41.9 69.7 90.2
[…]
Llama-3-8B-Instruct-262k 44.2 21.0 34.3 26.4 46.1 26.3 50.2 4.7 95.0 30.5 43.3 41.1 68.3 86.8
MInference 44.1 20.5 34.3 25.8 46.2 26.3 50.2 4.3 95.0 30.7 44.7 40.5 68.3 86.8
[…]
FlexPrefill; $\gamma=0.90$ 33.6 16.8 32.7 23.0 46.8 25.6 45.6 2.3 8.0 23.2 41.4 36.8 61.0 73.3
[…]
FlexPrefill; $\gamma=0.95$ 35.0 17.7 33.0 25.1 36.8 26.1 49.1 3.3 13.7 28.3 36.2 39.1 65.0 81.7
[…]
Qwen2.5-7B-Instruct 46.8 24.5 30.7 32.2 62.2 22.3 41.3 13.1 99.3 24.4 60.7 42.5 67.0 88.7
MInference 46.8 24.5 30.7 32.7 61.6 22.3 40.8 12.8 100.0 24.3 60.2 42.1 68.0 89.3
[…]
FlexPrefill; $\gamma=0.90$ 37.3 17.5 29.3 27.6 64.0 21.8 39.0 7.0 28.7 15.1 51.4 39.3 64.7 79.3
[…]
FlexPrefill; $\gamma=0.95$ 40.8 20.2 29.7 33.9 66.8 21.8 40.4 4.0 48.7 17.1 54.8 40.5 67.3 85.5
[…]
Table 8: Comparison with dynamic sparsity methods on LongBench.

### v2

@@
-TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling
+TriangleMix: Accelerating Prefilling via Decoding-time Contribution Sparsity
 Abstract
-Large Language Models (LLMs) rely on attention mechanisms whose time complexity grows quadratically with input sequence length, creating significant computational bottlenecks during the prefilling stage.
-Existing static sparse attention methods typically degrade accuracy, while dynamic sparsity methods introduce additional computational overhead due to runtime sparse index estimation.
-To address these limitations, we propose TriangleMix, a novel training-free static attention pattern.
-TriangleMix employs dense attention in shallow layers and switches to a triangle-shaped sparse pattern in deeper layers.
-Extensive experiments demonstrate that TriangleMix reduces attention overhead by 3.7×3.7$\times$3.7 × to 15.3×15.3$\times$15.3 × in deep layers, and decreases overall Time-to-First-Token (TTFT) by 12% to 32% for sequence lengths ranging from 32K to 128K, without sacrificing model accuracy.
-Moreover, TriangleMix can be seamlessly integrated with dynamic sparsity methods to achieve further speedup, e.g.
-accelerating MInference by 19% at 128K, highlighting its potential to enhance LLM inference efficiency.
-Implementation at https://aka.ms/TriangleMix
-TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling
-Zhiyuan He1, Yike Zhang2thanks: Work during internship at Microsoft. , Chengruidong Zhang1, Huiqiang Jiang1, Yuqing Yang1, Lili Qiu1 1Microsoft Research, 2Tsinghua University zhiyuhe@microsoft.com
+Large Language Models (LLMs) incur quadratic attention complexity with input length, creating a major time bottleneck in the prefilling stage.
+Existing acceleration methods largely exploit attention score sparsity by estimating blocks with high attention scores and applying dynamic sparse attention.
+In this work, we identify another untapped form of sparsity in the prefilling stage, namely decoding-time contribution sparsity, where many attention blocks exhibit nontrivial attention scores during prefilling yet contribute negligibly to subsequent decoding, as indicated by gradient-based analysis.
+Building on this observation, we propose TriangleMix, a training-free static attention pattern that uses dense attention in a subset of layers and switches to Triangle attention in the others.
+Extensive experiments show that TriangleMix preserves nearly lossless performance relative to dense attention while substantially reducing attention overhead in Triangle layers.
+For 128K inputs, Triangle attention achieves a $15.3\times$ speedup in attention computation, significantly exceeding the acceleration of typical dynamic sparse methods ($1.9\times$$3.4\times$).
+Furthermore, TriangleMix can be seamlessly combined with dynamic sparsity approaches, delivering an additional 6%–19% reduction in TTFT over using dynamic sparsity alone.
+(a) Attention sections.
+(b) Gradient of Middle Q-K with respect to the first generated token
+Figure 1: The average gradient $\mathrm{Grad}(\bm{M},l)$ of the Middle Q-K sections, measured on three models, shows a significant sparsity in deeper layers.
+This suggests that the Middle Q-K components in deeper layers contribute minimally to decoding and might potentially be skipped to improve efficiency.
+(a) Measure contribution based on different tokens
+(b) TriangleMix on Llama-3.1-8B-Instruct
+Figure 2: Left: Attention computation in certain layers exhibits contribution locality. Right: The proposed TriangleMix pattern for Llama-3.1-8B-Instruct.
 1 Introduction
-(a) TriangleMix
-(b) TriangleMix combined with dynamic attention
-Figure 1: TriangleMix on Llama-3.1-8B-Instruct.
-Large Language Models (LLMs) are capable of processing input sequences of varying lengths Grattafiori et al.
-(2024); Achiam et al.
-(2023); Yang et al.
-(2024), a crucial ability that supports diverse downstream tasks, such as question answering Bai et al.
-(2023), long-form document understanding Zhao et al.
-(2024), and code generation Jiang et al.
-(2024b).
-However, due to the quadratic complexity of attention mechanisms, the time-to-first-token (TTFT) significantly increases as the input context length grows.
-As prior research has demonstrated, attention computation has become a critical bottleneck in the prefilling stage of LLMs Jiang et al.
-(2024a); Lai et al.
-(2025).
-To address this bottleneck and accelerate the prefilling stage, researchers have proposed both static and dynamic sparse attention methods.
-Static sparse attention methods, such as StreamingLLM Xiao et al.
-(2023), reduce computational complexity from $O(N^{2})$ to $O(N)$ but suffer notable performance degradation on tasks with long contexts Li et al.
-(2024a).
-On the other hand, dynamic sparsity approaches, such as MInference Jiang et al.
-(2024a) and FlexPrefill Lai et al.
-(2025), aim to maintain near-full performance by dynamically computing sparse attention indices during inference.
-Dynamic sparsity methods are effective for handling extremely long contexts (e.g., beyond 128K tokens).
-However, they incur additional overhead due to the runtime estimation of sparse block indices.
-While this overhead is negligible at very long sequence lengths, it becomes non-trivial for moderately long contexts ranging from 32K to 128K tokens.
-We measured the average attention kernel time per layer on Llama-3.1-8B-Instruct, and the results (Table 1) show that MInference can be 2× slower than dense attention at 32K, and offers only limited speedup at 64K.
-Similarly, FlexPrefill delivers a modest 1.4× speedup at 32K.
-Figure 2: The average gradient $\mathrm{Grad}(\bm{M},l)$ of the Middle Q-K sections, measured on Llama-3.1-8B-Instruct, shows a significant decline in deeper layers.
-This suggests that the Middle Q-K components in deeper layers contribute minimally and might potentially be skipped to improve efficiency.
-In this work, we aim to develop a static attention pattern that effectively accelerates the prefilling stage across various context lengths.
-To achieve this goal, we propose a novel gradient-based method to identify the crucial sections within attention maps.
-Our analysis demonstrates that the middle Q-K section (Figure 2 and 3) of the attention map can be safely omitted in deeper layers without incurring performance loss.
-Motivated by this insight, we introduce TriangleMix, a lossless and efficient static attention pattern for accelerating long-context LLMs.
-As depicted in Figure 1, TriangleMix applies standard dense attention in shallow layers, and switches to a triangle-shaped sparse attention pattern in deeper layers.
-TriangleMix offers four major advantages:
-  • It is training-free, enabling direct integration with state-of-the-art pretrained LLMs.
-  • It achieves nearly lossless accuracy, preserving the full capabilities of the underlying models.
-  • Its static triangle-shaped sparsity eliminates the need for dynamic block index estimation and reduces the computational complexity from $O(N^{2})$ to $O(N)$.
-  • It is complementary to dynamic attention approaches.
-Replacing dynamic attention in deep layers with the static Triangle attention further enhances acceleration, compared with dynamic sparse attention alone.
-We conduct extensive experiments on three long-context LLMs, including Llama-3.1-8B-Instruct Grattafiori et al.
-(2024), Llama-3-8B-Instruct-262K GradientAI (2024), and Qwen2.5-7B-Instruct Yang et al.
-(2024), using all tasks from the RULER and LongBench benchmarks Hsieh et al.
-(2024); Bai et al.
-(2023).
-The results demonstrate that TriangleMix maintains the same accuracy as full attention.
-For Llama-3.1-8B-Instruct, TriangleMix decreases overall TTFT by 12% to 32% across context lengths ranging from 32K to 128K.
-Moreover, it can be seamlessly combined with dynamic sparsity methods without degrading performance.
-For example, integrating TriangleMix with MInference decreases TTFT by 19% at 128K context length for Llama-3.1-8B-Instruct.
+Large Language Models (LLMs) are capable of processing input sequences of varying lengths (Grattafiori et al., 2024; Achiam et al., 2023; Yang et al., 2024), a crucial ability that supports diverse downstream tasks, such as question answering (Bai et al., 2023), long-form document understanding (Zhao et al., 2024), and code generation (Jiang et al., 2024b).
+However, due to the quadratic complexity of attention mechanisms, the attention computation time significantly increases as the input context length grows.
+As prior research has demonstrated, attention computation has become a critical bottleneck in the prefilling stage of LLMs (Jiang et al., 2024a; Lai et al., 2025).
+To mitigate this bottleneck, several dynamic sparse attention techniques have been proposed, including MInference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and XAttention (Xu et al., 2025).
+These methods exploit attention score sparsity to accelerate computation: the majority of entries in attention matrix have negligible values and can be safely skipped, while blocks with significant scores need to be computed.
+Dynamic sparsity methods first estimate which regions are likely to contain high attention scores, and then selectively perform computation on these regions to approximate full attention (Jiang et al., 2024a; Lai et al., 2025; Xu et al., 2025).
+Dynamic sparsity attention is built on a simple intuition: low-score blocks can be skipped, while blocks with non-trivial scores must be computed.
+This raises a deeper question: Must we compute all non-trivial blocks during prefilling to preserve decoding accuracy?
+A natural concern is that skipping them might distort the attention distribution and degrade performance.
+However, our analysis reveals that the situation is subtler.
+We uncover a new form of sparsity in the prefilling stage, which we call decoding-time contribution sparsity.
+Unlike score sparsity, it reflects that certain blocks, though holding non-trivial scores, have little impact on the actual decoding process.
+To rigorously evaluate the necessity of each block, we analyze how its removal affects the generation of the first token following the prompt.
+Gradient analysis offers a direct measure of this sensitivity.
+As shown in Figure 1(a), we divide causal attention into three regions: Last Q–K, Middle Q–K, and the Streaming region.
+Our results indicate that the Streaming and Last Q–K regions are critical to decoding, while the Middle Q–K exhibits significant layer-wise sparsity pattern (Figure 1(b)).
+Notably, although the attention score magnitude of Middle Q–K regions are often comparable to Last Q–K, its gradients reveal substantially lower impact in certain layers.
+This sparsity arises from the phenomenon that certain parts of attention scores’ contribution has locality of influence with respect to the decoding token position.
+As shown in Figure 2(a), many attention scores only affect predictions within a limited temporal window after they occur.
+The Middle Q-K region in some layers has considerable influence on an intermediate token $o_{i}$, but contribute very little to the tokens generated after the entire prompt $o_{N},o_{N+1},$....
+This distinction allows us to safely skip Middle Q-K region’s attention computation of certain layers during the prefilling phase.
+Motivated by this insight, we introduce TriangleMix, a simple yet effective static attention pattern for accelerating prefilling.
+As depicted in Figure 2(b), TriangleMix combines dense attention in some layers with a triangle-shaped sparse attention pattern in others.
+This design brings four key benefits:
+  • Training-free: It can be applied directly to state-of-the-art pretrained LLMs without fine-tuning.
+  • Nearly lossless accuracy: Despite its simplicity, TriangleMix preserves model performance, reaching accuracy comparable to sophisticated dynamic sparsity methods.
+  • Efficiency: Its static triangle pattern removes the need for block index estimation, reducing complexity from $O(N^{2})$ to $O(N)$ and enabling significantly simpler and faster attention kernel than dynamic sparsity attention.
+  • Complementary to dynamic attention: Replacing dynamic sparsity with Triangle attention in selected layers delivers extra acceleration while maintaining performance.
+We conduct extensive experiments on three long-context LLMs, including Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Llama-3-8B-Instruct-262K (GradientAI, 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024), using all tasks from the RULER and LongBench benchmarks (Hsieh et al., 2024; Bai et al., 2023).
+Our results show that TriangleMix preserves the accuracy of full attention while substantially improving efficiency.
+Specifically, with 128K-token inputs, Triangle attention achieves a $15.3\times$ speedup in attention computation, far exceeding typical dynamic sparse methods ($1.9\times$$3.4\times$).
+For Llama-3.1-8B-Instruct, TriangleMix reduces overall TTFT by 12%–32% across context lengths from 32K to 128K.
+In addition, TriangleMix integrates seamlessly with dynamic sparsity approaches, yielding a further 6% to 19% decrease in TTFT compared to dynamic sparsity alone.
 2 Methodology
-2.1 Sparse Attention
-Figure 3: Three attention sections.
-Figure 4: First row: Average attention score $\mathrm{Att}(\bm{M},l)$ for the Middle and Last Q-K sections; Second row: average gradient $\mathrm{Grad}(\bm{M},l)$ for the Middle and Last Q-K sections.
-Method 32K 64K 128K
-Dense 44 179 750
-MInference 94 152 229
-FlexPrefill ($\gamma=0.95$) 32 71 220
-Table 1: Averaged attention kernel time in milliseconds for one layer on Llama-3.1-8B-Instruct.
+2.1 Probing Attention Block Contribution
 The prefilling stage of Transformer attention can be formulated as:
 $\bm{A}=\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M}))$
-[…]
-To accelerate computing, sparse attention aims to find a sparse mask matrix $\bm{M}^{\prime}$ to compute the attention output:
+where $\bm{Q},\bm{K},\bm{V}$ are matrices of shape $(N,d)$, and $\bm{M}$ is a causal mask matrix of shape $(N,N)$, with entries $\bm{M}_{i,j}\in\{0,1\}$.
+Here, $N$ represents the number of input tokens, and $c$ is a large positive constant to ensure attention scores masked by $\bm{M}_{i,j}=0$ become effectively zero after the softmax operation.
+To accelerate computing, sparse attention (Jiang et al., 2024a; Lai et al., 2025; Xu et al., 2025) aims to find a sparse mask matrix $\bm{M}^{\prime}$ to compute the attention output:
 $\bm{A^{\prime}}=\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M^{\prime}}))$
-where $|\bm{A}-\bm{A}^{\prime}|$ is expected to be sufficiently small to preserve the original model performance.
-Existing sparse attention methods fall into two categories: static sparsity and dynamic sparsity.
-In static sparsity, the sparse mask $\bm{M}^{\prime}$ is fixed and does not depend on the input.
-For example, StreamingLLM shows that most attention scores accumulate on the first few sink tokens and nearby tokens in a sliding window Xiao et al.
-(2023).
-This leads to the following streaming pattern:
-$\bm{M}^{\mathrm{streaming}}_{i,j}=\begin{cases}1,&i\geq j,j\leq si\\ 1,&i\geq j,i-j\leq sl\\ 0,&\mathrm{otherwise}\end{cases}$
-where $si$ is the number of sink tokens and $sl$ is the sliding window size.
-This pattern, also called "A-shape" in previous works Jiang et al.
-(2024a), is very efficient with only $O(N)$ complexity.
-However, it causes significant performance drops for long-context LLMs Li et al.
-(2024a).
-In contrast, dynamic sparsity methods, such as MInference Jiang et al.
-(2024a) and FlexPrefill Lai et al.
-(2025), estimate $\bm{M}^{\prime}$ at inference time based on the input through an observation window or the pooling vector of $\bm{Q}$ and $\bm{K}$.
-Dynamic sparsity generally provides better accuracy and acceleration for extremely long contexts (e.g., contexts exceeding 128K tokens).
-However, for moderately long contexts, the online estimation introduces additional time overhead.
-Specifically, as demonstrated in Table 1, MInference is slower than dense attention or offers small acceleration for contexts below 64K, while FlexPrefill achieves only modest improvements (approximately 1.4x speedup for 32K context).
-2.2 Attention Sections
-In this paper, we aim to design a static sparse attention mask that is both efficient and effective.
-Although a large fraction of attention scores falls into the streaming pattern, a significant performance gap still remains between streaming sparse attention and full dense attention.
-This leads us to the following question:
-Beyond the streaming pattern, what other parts of the attention are crucial for maintaining the end-to-end performance of long-context LLMs?
-To answer this, we divide the full attention matrix into three distinct sections as illustrated in Figure 3:
-  • Streaming section: includes attention sink and the sliding window;
-  • Last Q-K section: covers interactions between the last part of $\bm{Q}$ and $\bm{K}$, excluding the Streaming section;
-  • Middle Q-K section: consists of the remaining interactions between the middle parts of $\bm{Q}$ and $\bm{K}$.
-We define the Last Q-K mask as:
-$\bm{M}^{\mathrm{last}}_{i,j}=\begin{cases}1,&i\geq j,N-i<last,\\ &j>si,i-j>sl\\ 0,&\mathrm{otherwise}\end{cases}$
-[…]
-Similarly, the Middle Q-K mask is defined as:
-$\bm{M}^{\mathrm{middle}}_{i,j}=\begin{cases}1,&i\geq j,N-i\geq last,\\ &j>si,i-j>sl\\ 0,&\mathrm{otherwise}\end{cases}$
-2.3 Probing Attention Importance
-Firstly, we inspect the average attention scores within the two sections defined by masks $\bm{M}^{\mathrm{middle}}$ and $\bm{M}^{\mathrm{last}}$.
-We observe that these average attention scores are quite similar, although both are notably lower than the average score observed in the streaming section $\bm{M}^{\mathrm{streaming}}$.
-To achieve a more precise measurement of the relative importance of these sections, we introduce a novel gradient-based probing technique.
-Specifically, given an input $\bm{X}_{\mathrm{input}}$ containing $N$ tokens fed into a language model, we define a probing variable $\bm{\theta}\in\mathbb{R}^{L\times N\times N}$, where $\bm{\theta}_{\ell,i,j}=1$ for all layers $\ell$ and token indices $i,j$.
+The mask matrix $\bm{M^{\prime}}$ can be either static or dynamic.
+StreamingLLM (Xiao et al., 2023) and LM-infinite (Han et al., 2024) employ similar static streaming mask, which restricts attention to a few sink tokens and a sliding window of nearby tokens.
+Such pattern reduces the computation complexity of attention to $O(N)$ but significantly harms the performance of long-context LLMs(Li et al., 2024a).
+In contrast, dynamic masks enable dynamic sparsity attention.
+During inference, block indices in $\bm{M^{\prime}}$ that likely yield non-trivial scores are identified from the input, and FlashAttention (Dao et al., 2022) is applied only to these blocks.
+Methods such as MInference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and XAttention (Xu et al., 2025) all adopt this strategy.
+A central challenge for these dynamic sparsity attention lies in efficiently and accurately estimating the block indices with meaningful attention scores.
+MInference categorizes attention heads into three types and applies different estimation strategies, either by observing recent query tokens or by pooling vectors from $\bm{Q}$ and $\bm{K}$.
+FlexPrefill adopts a similar strategy but determines the head type dynamically during inference.
+XAttention, on the other hand, uses the sum of anti-diagonal scores to identify blocks with large attention.
+Despite these differences, the common objective remains the same: reliably locating blocks with meaningful attention scores.
+In this paper, we move beyond block index estimation and pose a more fundamental question: Must we compute all non-trivial blocks during prefilling to preserve decoding accuracy?
+To answer this, we quantitatively probe each block’s contribution relative to the generated tokens in decoding, which we call its decoding-time contribution.
+Our goal is to prioritize computation for blocks with higher decoding-time contribution.
+Since gradients naturally capture such sensitivity, we propose a gradient-based method to perform this analysis.
+Given an input $\bm{X}_{\mathrm{input}}$ containing $N$ tokens fed into a language model, we define a probing variable $\bm{\theta}\in\mathbb{R}^{L\times N\times N}$, where $\bm{\theta}_{\ell,i,j}=1$ for all layers $\ell$ and token indices $i,j$.
 Here, $L$ represents the total number of layers in the model.
-For simplicity of notation, we omit the attention head dimension here, but our proposed method generalizes naturally to both Multi-head Attention Vaswani et al.
-(2017) and Grouped Query Attention Ainslie et al.
-(2023).
+For simplicity of notation, we omit the attention head dimension here, but our proposed method generalizes naturally to both Multi-head Attention (Vaswani et al., 2017) and Grouped Query Attention (Ainslie et al., 2023).
 In layer $l$, the attention scores is calculated as:
@@
 where $\odot$ denotes element-wise multiplication. Since all elements of $\bm{\theta}$ are initially set to 1, this operation does not alter the attention scores.
-The model then outputs a logit prediction $\bm{Y_{\theta}}$ for the next token, represented as a vector $\bm{Y_{\theta}}=[y_{1},y_{2},\dots,y_{\text{vocab}}]$.
+The model then outputs a logit prediction $\bm{Y_{\theta}}$ for the next token after the $N$th token, represented as a vector $\bm{Y_{\theta}}=[y_{1},y_{2},\dots,y_{\text{vocab}}]$.
+Note that it is only the output of the last query.
 Our focus is on a specific logit $y_{\mathrm{gt}}$, associated with the correct label or ground truth token.
@@
 We are particularly interested in the partial derivative $\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}$, which measures the sensitivity of the model’s output to changes in attention scores.
-We quantify the importance of a specific attention section by computing the mean of the derivative values within that section.
+We quantify the importance of a specific attention section $\bm{M}^{\prime}$ by computing the mean of the derivative values within that section.
 Formally, we define the gradient-based importance as:
-$\mathrm{Grad}(\bm{M},l)=\frac{\sum_{\bm{M}_{i,j}=1}{\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}}}{\sum_{\bm{M}_{i,j}=1}{\bm{M}_{i,j}}}$
-where $\bm{M}$ denotes the binary mask representing the targeted attention section, and $\mathrm{Grad}(\bm{M},l)$ indicates the average sensitivity of $y_{\mathrm{gt}}$ to changes within this section at layer $l$.
-For comparison, we also define the average attention score of the section as:
-$\mathrm{Att}(\bm{M},l)=\frac{\sum_{\bm{M}_{i,j}=1}{\bm{A}_{l,i,j}}}{\sum_{\bm{M}_{i,j}=1}{\bm{M}_{i,j}}}$
+$\mathrm{Grad}(\bm{\hat{M}},l)=\frac{\sum_{\bm{\hat{M}}_{i,j}=1}{\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}}}{\sum_{\bm{\hat{M}}_{i,j}=1}{\bm{\hat{M}}_{i,j}}}$
 […]
-Empirically, we leverage a key-value retrieval task to analyze the importance of different sections of attention.
-The language model is tasked with retrieve the corresponding value, starting from its first generated token.
-$y_{gt}$ is set to be the first token of the correct value.
-The full task prompt is provided in Appendix A.1.
-We calculate $\mathrm{Grad}(\bm{M},l)$ and $\mathrm{Att}(\bm{M},l)$ for both the middle Q-K section ($\bm{M}^{\mathrm{middle}}$) and the last Q-K section ($\bm{M}^{\mathrm{last}}$) across 100 randomly generated samples.
-The hyperparameters are set as follows: the number of sink tokens ($si$) is 64, the sliding window size ($sl$) is 128, and the length of the last attention section ($last$) is also set to 128.
-Each input sequence consists of approximately 2,000 tokens.
-Finally, we visualize the results for $\mathrm{Grad}(\bm{M},l)$ and $\mathrm{Att}(\bm{M},l)$ in Figure 4.
-For clarity, attention scores and gradient values are both summed across all heads within each layer.
+2.2 Decoding-time Sparsity in Middle Q-K sections
+Figure 3: First row: Average attention score for the Middle and Last Q-K sections; Second row: average gradient $\mathrm{Grad}(\bm{M},l)$ for the Middle and Last Q-K sections.
+Under the framework proposed in Section 2.1, we conduct an initial gradient-based analysis of three predefined attention sections.
+As illustrated in Figure 1(a), we partition attention into the Streaming section, the Middle Q–K section, and the Last Q–K section.
+We evaluate these regions using a key–value retrieval task in which the language model must output the corresponding value beginning with its first generated token; we set $y_{\mathrm{gt}}$ to the first token of the correct value.
+See Appendix A.1 for the exact region definitions and the full prompt.
+First, we observe that both the attention scores and the gradients are high in the Streaming section, consistent with prior works (Xiao et al., 2023; Han et al., 2024).
+For the Middle Q–K and Last Q–K sections, we summarize our findings as follows:
+Slightly higher attention scores in Middle Q–K.
+As shown in the top row of Figure 3, the average attention scores in the Middle Q–K and Last Q–K sections are similar, with the Middle Q–K section being slightly higher.
+This indicates that both regions contain non-trivial attention blocks.
+Lower gradients in Middle Q–K.
+Despite the slightly higher attention scores, the Middle Q–K section exhibits substantially smaller gradients than the Last Q–K section (The bottom row of Figure 3).
+This suggests that, for predicting the first token, computing blocks in the Last Q–K region is more critical.
+Layer-wise sparsity.
+The gradients within the Middle Q–K section display clear layer-wise sparsity.
+For example, in Llama-3.1-8B-Instruct and Llama-3-8B-262K (Figure 1(b)), the gradients of the last 16 layers are near zero, implying that block computations in this region could be skipped with minimal impact on first-token prediction.
+We refer to this phenomenon as decoding-time contribution sparsity, since it is defined directly by contributions to decoded tokens.
+Moreover, the observation extends beyond the first token: for later generated tokens, the current Middle Q–K region becomes a subset of the future Middle Q–K region.
+2.3 Perplexity Analysis
+Llama-3.1-8B-Instruct Llama-3-Instruct-262k Qwen2.5-7B-Instruct
+Method 1024–1152 1920–2048 1024–1152 1920–2048 1024–1152 1920–2048
+Dense 8.31 7.81 8.05 7.62 8.05 7.62
+Skip Q (512–1024)-K 50% layers 8.58 (+0.27) 7.82 (+0.01) 8.33 (+0.28) 7.63 (+0.01) 8.16 (+0.11) 7.64 (+0.02)
+Skip Q (512–1024)-K all layers 8.76 (+0.45) 7.90 (+0.09) 8.54 (+0.49) 7.72 (+0.10) 8.54 (+0.49) 7.72 (+0.10)
+Table 1: Perplexity (PPL) comparison across different models and settings.
+To further validate our claim, we study the impact of removing attention computation in the middle region.
+We randomly sample 1,024 Wikipedia articles and take the first 2,048 tokens from each.
+Three settings are compared: (1) full attention, (2) skipping attention with Q indices from 512 to 1,024 across all layers, and (3) skipping the same region in only half of the layers.
+In both (2) and (3), the Streaming section and the attention with other Q indices are still retained.
+For the third setting, we select the half of layers with the lowest $\mathrm{Grad}(\bm{\hat{M}},l)$, where $\bm{\hat{M}}$ denotes the skipped attention region and the gradient measures its contribution to predicting the 2049th token.
+For each model, the skipped layers remain fixed across all 1,024 samples, although different models may choose different sets of layers.
+Table 1 reports perplexity changes for token ranges 1024–1152 and 1920–2048.
+At position 1024, the skipped computation corresponds to the Last Q–K section (since Q indices 512–1024 are omitted).
+At positions 1920–2048, the skipped region aligns with the Middle Q–K section.
+We find that skipping either all or half of the layers significantly increases perplexity in the 1024–1152 range, confirming our observations in Section 2.2 that the Last Q–K section is critical.
+In contrast, at 1920–2048, skipping half of the layers causes only a minor increase in perplexity (0.01–0.02), though skipping all Middle Q–K computations still raises perplexity noticeably.
+We visualize this phenomenon in Figure 2(a).
+We refer to it as the locality of influence, which describes the finite-horizon effect of certain attention scores in transformer models.
+Specifically, even if some attention blocks have non-trivial scores, their influence on the output distribution disappears once the decoding position $n$ exceeds $t+\Delta t$, where $t$ is the query index of the block and $\Delta t$ is a small constant (e.g., $\Delta t=128$).
+This indicates that many attention scores only affect predictions within a limited temporal window after they occur.
+As a result, they influence nearby predictions but do not affect distant tokens, particularly those beyond the prompt.
+This distinction allows us to safely skip computing certain attention weights during the prefilling phase without relying on the magnitude of the attention weights themselves.
 2.4 TriangleMix
+Based on these observations, we introduce TriangleMix, an effective and efficient static attention pattern for long-context prefilling.
+The key idea is that the Middle Q–K section in some layers contributes little to decoding, and can therefore be safely skipped.
+In these layers, attention reduces to the Triangle pattern shown in Figure 2(b).
+This modification lowers the complexity of attention in those layers from $O(N^{2})$ to $O(N)$.
+Unlike dynamic sparsity, Triangle attention is static: there is no need to predict block indices or design specialized sparse kernels.
+As a result, the kernel is much simpler and faster to implement.
+Overall, TriangleMix combines dense attention in some layers with Triangle attention in others.
+For a given model, the application of TriangleMix is as follows.
+We first conduct the gradient-based analysis introduced in Sections 2.1 and 2.2.
+For each layer $l_{i}$, we compute the decoding-time contribution of its Middle Q–K section, denoted as $\mathrm{Grad}(\bm{M}^{\mathrm{middle}},l_{i})$ for $i=1,2,\dots,N_{\mathrm{layer}}$.
+Based on these values, we identify the $L_{\mathrm{tri}}$ layers with the lowest contributions, where $L_{\mathrm{tri}}$ is a hyperparameter controlling how many layers are converted into Triangle attention.
+These selected layers adopt the Triangle pattern at inference time, thereby reducing computation.
+The remaining layers retain their original dense attention.
+Alternatively, they can also be replaced with dynamic sparse attention methods to achieve further acceleration.
+3 Evaluations
+3.1 Settings
+LLMs and Benchmarks.
+We evaluate our method on three recent long-context LLMs: Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Llama-3-8B-Instruct-262K (GradientAI, 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024).
+We use two challenging benchmarks: LongBench (Bai et al., 2023) and RULER (Hsieh et al., 2024).
+On RULER, we test input lengths of 4K, 8K, 16K, 32K, 64K, and 128K.
+On LongBench, we evaluate all English tasks.
+Hyperparameters.
+Different models exhibit varying levels of decoding-time contribution sparsity.
+We set $L_{\mathrm{tri}}=16$ for Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, and $L_{\mathrm{tri}}=8$ for Qwen2.5-7B-Instruct, which preserves nearly lossless accuracy.
+An analysis of different $L_{\mathrm{tri}}$ settings is provided in Section 4.
+For all experiments, we fix the sink token size to 8 and the local window size to 512.
+Static Sparsity Baselines.
+We compare our approach with four static sparsity baselines: (1) Streaming pattern (Xiao et al., 2023; Han et al., 2024), applied to all layers during prefilling; (2) Triangle attention, also applied to all layers; (3) StreamingMix, where Triangle attention in TriangleMix is replaced by the Streaming pattern; (4) DuoAttention (Xiao et al., 2024), which learns a separate sparsity pattern for each attention head.
+For DuoAttention, we use the official pattern for Llama-3.1-8B-Instruct and train a pattern for Llama-3-8B-Instruct-262K using the authors’ script.
+We set the sparsity ratio to be 50%.
+However, training on Qwen2.5-7B-Instruct did not converge, so results are omitted.
+For all static sparsity baselines, we set the sink token number and local window to 8 and 512, consistent with our method.
+Methods 4K 8K 16K 32K 64K 128K Avg.
+Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
+Streaming 64.1 55.4 40.5 28.9 26.7 3.3 36.5
+Triangle 88.8 88.3 82.8 72.6 65.0 39.0 72.7
+StreamingMix 94.3 91.8 90.2 86.2 79.2 71.9 85.6
+DuoAttention 95.7 93.1 88.4 84.3 82.5 64.0 84.7
+TriangleMix 96.3 95.1 94.7 91.3 86.3 77.5 90.2
+Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
+Streaming 49.4 38.7 33.4 30.2 26.5 21.7 33.3
+Triangle 88.1 84.7 80.6 71.0 65.0 55.4 74.1
+StreamingMix 87.3 83.2 80.8 79.9 77.6 70.8 79.9
+DuoAttention 92.8 91.9 89.0 85.0 81.1 76.1 86.0
+TriangleMix 93.5 91.0 88.1 85.0 82.4 79.6 86.6
+Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
+Streaming 55.6 47.8 35.9 30.6 28.1 21.0 36.5
+Triangle 88.4 81.9 74.0 58.3 51.0 14.6 61.4
+StreamingMix 93.3 90.4 89.4 81.7 76.4 63.6 82.5
+TriangleMix 95.5 93.8 92.1 84.6 80.2 67.7 85.6
+Table 2: Comparison with static sparsity methods on RULER.
+Llama-3.1, Llama-3-262K, Qwen2.5 are abbreviations for Llama-3.1-8B-Instruct, Llama-3-8B-262K, and Qwen2.5-8B-Instruct.
+The same applies to Table 3.
+Methods 4K 8K 16K 32K 64K 128K Avg.
+Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
+TriangleMix 96.3 95.1 94.7 91.3 86.3 77.5 90.2
+MInference 96.3 95.1 95.0 90.5 86.8 75.0 89.8
+Ours + MInfer 96.2 95.0 94.5 90.8 87.2 75.8 89.9
+FlexPrefill 95.0 94.7 94.5 92.8 86.5 75.9 89.9
+Ours + FP 95.2 95.0 94.7 92.8 86.8 76.2 90.1
+XAttn 96.5 94.9 94.5 91.8 85.7 73.8 89.5
+Ours + XAttn 96.2 95.1 94.9 91.9 85.7 72.9 89.4
+Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
+TriangleMix 93.5 91.0 88.1 85.0 82.4 79.6 86.6
+MInference 93.4 90.7 88.8 85.1 82.6 80.1 86.8
+Ours + MInfer 93.0 90.4 88.6 84.4 82.7 80.1 86.5
+FlexPrefill 90.3 87.7 88.0 83.8 80.2 75.7 84.3
+Ours + FP 90.2 87.3 87.3 83.8 80.0 76.6 84.2
+XAttn 93.6 91.1 87.7 84.8 82.6 78.9 86.5
+Ours + XAttn 93.3 91.2 87.9 84.3 82.5 79.3 86.4
+Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
+TriangleMix 95.5 93.8 92.1 84.6 80.2 67.7 85.6
+MInference 95.6 93.8 92.8 84.9 79.1 68.0 85.7
+Ours + MInfer 95.7 93.5 92.4 84.8 76.9 63.5 84.5
+FlexPrefill 91.1 89.7 88.5 75.4 72.5 51.2 78.1
+Ours + FP 92.2 91.1 89.5 77.0 73.3 52.7 79.3
+XAttn 95.4 93.7 92.0 82.9 77.8 66.2 84.7
+Ours + XAttn 95.3 93.3 91.5 82.2 77.4 64.9 84.1
+Table 3: Comparison with dynamic sparsity methods on RULER.
+Dynamic Sparsity Baselines.
+We further compare with three dynamic sparsity methods: MInference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and XAttention (Xu et al., 2025).
+Our method can also be combined with these approaches: Triangle attention is retained, while layers that require full attention are replaced with dynamic sparse attention.
+This design further reduces runtime overhead compared with dynamic sparsity alone, since Triangle attention skips more attention computation and is much simpler.
+We denote the combined versions as “Ours + MInference,” “Ours + FlexPrefill,” and “Ours + XAttention.” For dynamic methods, we also set their hyperparameters to maintain near-lossless accuracy: $\gamma=0.95$ for FlexPrefill, and $\tau=0.95$ with $\mathrm{stride}=8$ for XAttention.
+3.2 Effectiveness of TriangleMix
+Comparison with Static Sparsity.
+Table 3 and Table 5 present the evaluation results of TriangleMix and various static sparsity baselines.
+TriangleMix consistently outperforms all static baselines and remains nearly lossless performance across all input lengths.
+In contrast, other static methods usually suffer noticeable performance degradation.
+DuoAttention performs comparably to TriangleMix on LongBench using Llama-3.1-8B-Instruct.
+However, its performance is worse than TriangleMix on RULER, especially deteriorates at longer lengths (e.g., 64K and 128K).
+Furthermore, DuoAttention requires a separate training phase to learn head-wise sparsity, which introduces additional computational overhead and may fail to converge on models such as Qwen-2.5-7B-Instruct.
+On the LongBench benchmark, we observe that the full Triangle pattern performs poorly on PassageRetrieval tasks, while the StreamingMix pattern underperforms on the in-context learning task (TREC).
+These findings suggest that attention over the Middle Q-K section in certain layers (especially shallow layers) is still crucial for retrieval tasks, while attention over the Last Q-K section in certain layers (especially deep layers) plays a key role in supporting in-context learning.
+Methods Average
+2Wiki
+GovRep
+Hotpot
+LCC
+MNews
+MF-en
+PsgCnt
+PsgRtr
+Qasper
+Rbench
+Samsum
+TREC
+TrvQA
+Llama-3.1-8B-Instruct 54.8 48.1 34.4 61.6 66.6 25.6 56.9 16.8 99.7 44.9 52.7 42.6 71.0 91.6
+TriangleMix 54.5 47.8 34.3 61.5 67.0 25.8 56.5 16.8 98.3 44.8 51.9 42.2 69.7 92.2
+MInference 54.9 48.1 34.5 61.4 67.0 25.9 57.0 17.5 99.7 44.7 52.5 42.6 71.3 91.9
+Ours + MInference 54.5 48.3 34.2 62.0 66.7 25.6 56.6 17.4 98.3 44.8 51.6 42.1 69.3 92.1
+FlexPrefill 48.4 39.4 33.7 58.3 67.2 25.6 55.6 4.0 47.7 42.5 53.8 41.9 69.7 90.2
+Ours + FlexPrefill 52.2 39.4 33.5 58.5 66.9 25.7 54.5 3.3 97.0 43.1 53.8 42.3 69.3 90.8
+XAttention 53.5 45.4 34.5 61.0 66.9 25.7 55.9 18.1 86.0 42.9 55.5 42.9 71.0 90.2
+Ours + XAttention 54.3 45.0 34.6 61.0 66.8 25.6 57.1 19.0 98.7 42.4 54.4 42.0 69.3 90.4
+Llama-3-8B-Instruct-262k 44.2 21.0 34.3 26.4 46.1 26.3 50.2 4.7 95.0 30.5 43.3 41.1 68.3 86.8
+TriangleMix 43.4 20.7 34.5 25.7 47.3 26.1 51.1 4.7 85.3 30.7 43.6 40.9 66.7 87.0
+MInference 44.1 20.5 34.3 25.8 46.2 26.3 50.2 4.3 95.0 30.7 44.7 40.5 68.3 86.8
+Ours + MInference 43.4 20.6 34.4 26.4 47.2 26.2 51.1 4.3 85.3 30.3 43.6 40.9 66.7 86.9
+FlexPrefill 35.0 17.7 33.0 25.1 36.8 26.1 49.1 3.3 13.7 28.3 36.2 39.1 65.0 81.7
+Ours + FlexPrefill 36.0 18.3 33.2 25.1 36.9 26.2 49.6 6.7 24.0 28.3 35.6 38.9 64.3 80.4
+XAttention 41.5 20.7 34.3 26.0 37.9 26.1 52.3 4.7 71.3 30.8 39.4 40.0 68.3 87.2
+Ours + XAttention 42.1 21.1 34.3 26.1 37.7 26.2 51.7 5.7 81.3 30.3 38.3 39.9 67.3 87.0
+Qwen2.5-7B-Instruct 46.8 24.5 30.7 32.2 62.2 22.3 41.3 13.1 99.3 24.4 60.7 42.5 67.0 88.7
+TriangleMix 46.1 24.7 30.6 32.3 62.6 22.3 40.9 13.1 89.3 23.9 59.6 42.4 69.3 88.5
+MInference 46.8 24.5 30.7 32.7 61.6 22.3 40.8 12.8 100.0 24.3 60.2 42.1 68.0 89.3
+Ours + MInference 46.1 25.7 30.3 32.2 61.9 22.3 40.7 11.4 89.7 25.8 59.4 42.5 68.7 89.0
+FlexPrefill 40.8 20.2 29.7 33.9 66.8 21.8 40.4 4.0 48.7 17.1 54.8 40.5 67.3 85.5
+Ours + FlexPrefill 41.7 21.0 29.8 32.8 64.8 21.9 39.6 5.3 56.0 19.8 55.1 40.9 69.0 85.8
+XAttention 46.6 22.4 30.9 34.0 63.9 22.4 40.5 10.9 98.0 23.5 60.9 42.2 67.7 88.8
+Ours + XAttention 45.8 23.6 31.1 33.1 63.8 21.7 41.3 9.8 88.3 24.5 58.2 42.4 69.0 88.4
+Table 4: Comparison with dynamic sparsity methods on LongBench.
+Comparison with Dynamic Sparsity.
+Tables 3 and 4 present the results of TriangleMix compared with dynamic sparsity methods, as well as their combinations.
+Despite its simplicity, TriangleMix maintains accuracy comparable to dynamic sparsity methods.
+On the RULER benchmark, it yields an average accuracy change of only $0.4\%$ to $0.1\%$ relative to full attention, which is on par with MInference ($0.6\%$ to $0.3\%$), and outperforms FlexPrefill ($9.1\%$ to $0.6\%$) and XAttention ($2.2\%$ to $0.0\%$).
+Importantly, although TriangleMix skips many attention blocks with non-trivial scores, the decoding accuracy remains nearly lossless.
+This supports our claim that such blocks contribute only marginally to decoding-time accuracy, and that decoding-time sparsity is a general property across all three tested models.
+Moreover, when combined with dynamic sparsity methods such as MInference, FlexPrefill, or XAttention, TriangleMix consistently preserves their original performance.
+On LongBench, combining TriangleMix with FlexPrefill or XAttention even yields slight performance improvements.
+These combinations are meaningful: by eliminating computation in many additional attention blocks, TriangleMix reduces the computation required by dynamic sparsity methods, making prefilling even faster than using dynamic sparsity alone.
+Overall, TriangleMix preserves nearly full accuracy of dense attention across tasks while significantly reducing time complexity from $O(N^{2})$ to $O(N)$ in layers using Triangle attention.
 Methods Average
@@
 DuoAttention 54.5 44.6 34.1 59.1 68.0 25.3 55.0 15.7 99.7 43.9 58.8 42.7 71.3 91.0
-[…]
+TriangleMix 54.5 47.8 34.3 61.5 67.0 25.8 56.5 16.8 98.3 44.8 51.9 42.2 69.7 92.2
 Llama-3-8B-Instruct-262k 44.2 21.0 34.3 26.4 46.1 26.3 50.2 4.7 95.0 30.5 43.3 41.1 68.3 86.8
@@
 DuoAttention 42.6 19.2 34.6 25.1 44.4 26.1 51.8 2.3 87.7 29.6 41.1 40.4 65.7 85.9
-[…]
+TriangleMix 43.4 20.7 34.5 25.7 47.3 26.1 51.1 4.7 85.3 30.7 43.6 40.9 66.7 87.0
 Qwen2.5-7B-Instruct 46.8 24.5 30.7 32.2 62.2 22.3 41.3 13.1 99.3 24.4 60.7 42.5 67.0 88.7
@@
 StreamingMix 45.2 24.4 31.9 38.5 68.2 22.5 43.5 10.7 93.0 24.8 60.9 42.0 43.7 83.1
-[…]
+TriangleMix 46.1 24.7 30.6 32.3 62.6 22.3 40.9 13.1 89.3 23.9 59.6 42.4 69.3 88.5
 Table 5: Comparison with static sparsity methods on Longbench.
-TriangleMix is our proposed method.
-StreamingMix is a hybrid baseline that uses dense attention in shallow layers but replaces Triangle attention with a streaming pattern in deeper layers.
-The number of dense layers matches that of TriangleMix.
-From Figure 4, we observe that both the Middle Q-K and Last Q-K sections have low average attention scores, with the Middle Q-K section slightly higher across all layers.
-However, when examining the gradient values, we find that the derivatives in the Last Q-K section are significantly larger than those in the Middle Q-K section.
-This indicates that the model output is more sensitive to changes in the attention scores of the Last Q-K section compared to the Middle Q-K section.
-Besides, for the last layer, we observe that we always have $\mathrm{Grad}(\bm{M},l)=0$.
-This phenomenon is expected due to the specific structure of the attention mechanism.
-In the last layer, the logit corresponding to the first predicted token is computed solely based on the last query vector and the entire key matrix $\bm{K}^{T}$.
-Formally, this corresponds to computing only the last row of the attention map, represented as $\mathrm{Softmax}(\bm{Q}_{\mathrm{last}}\bm{K}^{T})$.
-Since the gradient measure $\mathrm{Grad}(\bm{M},l)$ is defined based on the derivative of the ground truth output logit with respect to parameters associated with the attention sections defined by the mask, and because the first output token logit does not depend on the Middle Q-K sections in the last layer (which involve query vectors from other positions), the resulting gradient naturally becomes zero.
-During the training phase, this observation does not apply.
-The training loss is typically computed using all output logits across the entire input sequence, which requires calculating the full attention matrix in the last layer.
-However, during inference, since our primary interest lies only in the prediction, we can optimize the computation in the final layer by restricting attention calculation to $\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{Q}_{\mathrm{last}}\bm{K}^{T})$.
-Moreover, we observe that the gradient values for the Middle Q-K sections follow a consistent trend across all models: they remain moderate in the shallow layers but drop sharply in deeper layers.
-As shown in Figure 2, this pattern is evident.
-For Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, the decline occurs between layers 10 and 20, while for Qwen2.5-7B-Instruct, it starts after layer 20.
-This suggests that the attention from the Middle Q-K sections in deep layers contribute less to the output logits.
-Notably, the number of elements in the Middle Q-K sections scales quadratically with the input token length $N$, whereas the elements in the Streaming and Last Q-K sections grow only linearly with $N$.
-Therefore, if the computation of the Middle Q-K sections is skipped in deep layers, the time complexity of attention can be reduced from $O(N^{2})$ to $O(N)$.
-And we assume such skipping will not cause significant loss in output quality.
-Based on these observations, we propose TriangleMix, an effective and efficient static attention pattern designed for long-context LLMs.
-We denote the starting layer of Triangle attention as $L_{\mathrm{tri\_start}}$.
-For layer $l\leq L_{\mathrm{tri\_start}}$, we apply standard dense attention:
-$\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M})\right)$
-where $\bm{M}$ is the standard causal attention mask.
-In layer $l>L_{\mathrm{tri\_start}}$, we skip the Middle Q-K sections and apply a Triangle-shape attention pattern:
-$\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c\left(1-(\bm{M}-\bm{M}^{\text{middle}})\right)\right)$
-where $\bm{M}^{\text{middle}}$ represents the mask corresponding to the Middle Q-K sections.
-Thus, in deep layers, the attention complexity becomes $O(N)$, and there is no need for dynamic estimation of the attention mask, making the computation highly efficient.
-Furthermore, we can optionally adopt dynamic sparse attention in shallow layers to further accelerate computation, while keeping the static Triangle pattern in deep layers.
-We acknowledge that the Triangle-shape attention pattern was first introduced in Scbench Li et al.
-(2024a).
-However, Scbench applies the Triangle-shape attention uniformly across all layers, leading to notable performance degradation.
-In contrast, by using a gradient-based probing method, we find that the Middle Q-K sections remain important in shallow layers.
-Therefore, by combining dense (or dynamic sparse) attention in shallow layers and Triangle-shape attention in deep layers, TriangleMix preserves model performance while achieving significant efficiency gains.
-3 Evaluations
-3.1 Settings
-LLMs.
-We evaluate our method on the following long-context language models.
-(1) Llama-3.1-8B-Instruct, a state-of-the-art model from Meta with a context window of up to 128K tokens Grattafiori et al.
-(2024).
-(2) Llama-3-8B-Instruct-262K, a fine-tuned variant from Llama-3-8B-Instruct that supports up to 262K tokens GradientAI (2024).
-(3) Qwen2.5-7B-Instruct, a model from Alibaba with a maximum context window of 128K Yang et al.
-(2024).
-Static Sparsity Baselines.
-We compare our method with four static sparsity baselines.
-The first is the Streaming pattern Xiao et al.
-(2023), applied to all layers with $si=8$ and $sl=512$.
-The second is Triangle Attention applied to all layers, using $si=8$, $sl=512$, and $last=128$.
-The third is a hybrid baseline StremingMix, where dense attention is used for layers $l\leq L_{\text{tri\_start}}$, and streaming attention is used for deeper layers.
-We also compare against DuoAttention, which learns a separate sparsity pattern for each attention head Xiao et al.
-(2024).
-DuoAttention provides a trained pattern for Llama-3.1-8B-Instruct; we train a sparsity pattern for Llama-3-8B-Instruct-262K using the authors’ script.
-However, for Qwen2.5-7B-Instruct, the training does not converge, so its result is omitted.
-At the prefilling stage, half of the heads follow the streaming pattern, while the other half use dense attention.
-Dynamic Sparsity Baselines.
-We compare with two dynamic sparsity methods: MInference Jiang et al.
-(2024a) and FlexPrefill Lai et al.
-(2025).
-We apply default sparsity settings of all LLMs from MInference.
-For FlexPrefill, we test its hyperparameter $\gamma=0.90$ and $\gamma=0.95$.
-Additionally, our method can be combined with dynamic sparsity attention: Dynamic sparsity (MInference or FlexPrefill) is applied for layers $l\leq L_{\text{tri\_start}}$, while Triangle Attention is used for deeper layers.
-This combination reduces runtime overhead because Triangle Attention avoids runtime mask estimation in dynamic sparse attention.
-We use “Ours + MInference” and “Ours + FlexPrefill” to denote the corresponding combined versions.
-Hyperparameters.
-The hyperparameter $L_{\text{tri\_start}}$ determines the layer from which the Triangle Attention pattern is applied.
-For layers $l\leq L_{\text{tri\_start}}$, we use dense attention or dynamic sparsity methods.
-For layers $l>L_{\text{tri\_start}}$, we switch to the static Triangle Attention.
-Based on the observations in Figure 4, we set $L_{\text{tri\_start}}=16$ for Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, and $L_{\text{tri\_start}}=20$ for Qwen2.5-7B-Instruct.
-For the Triangle mask, we use: sink tokens $si=8$, sliding window size $sl=512$, and last window size $last=128$, which is the same with the sparse sparsity baselines.
-Benchmarks.
-We evaluate on two benchmarks: (1) RULER: A synthetic long-context benchmark including tasks like hay-in-the-stack, KV retrieval, variable tracking, and QA.
-We evaluate at input lengths of 4K, 8K, 16K, 32K, 64K, and 128K, with 100 samples per task-length pair Hsieh et al.
-(2024).
-(2) LongBench: A realistic long-context benchmark with tasks such as single-doc and multi-doc QA, few-shot learning, code generation, retrieval, summarization, and counting Bai et al.
-(2023).
-We evaluate all English tasks.
-Specifically, Qwen2.5-7B-Instruct adopts Yarn scaling to improve its long context ability for input length larger than 32K Peng et al.
-(2023).
-We report the performance of the origin Qwen2.5-7B-Instruct model for RULER length < 32K, and the Yarn-scaled model for RULER length 32K, 64K and 128K.
-For LongBench, we only evaluate the origin Qwen2.5-7B-Instruct model for simplicity.
-3.2 Effectiveness of TriangleMix
-Methods 4K 8K 16K 32K 64K 128K Avg.
-Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
-Streaming 64.1 55.4 40.5 28.9 26.7 3.3 36.5
-Triangle 88.8 88.3 82.8 72.6 65.0 39.0 72.7
-StreamingMix 94.3 91.8 90.2 86.2 79.2 71.9 85.6
-DuoAttention 95.7 93.1 88.4 84.3 82.5 64.0 84.7
-[…]
-Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
-Streaming 49.4 38.7 33.4 30.2 26.5 21.7 33.3
-Triangle 88.1 84.7 80.6 71.0 65.0 55.4 74.1
-StreamingMix 87.3 83.2 80.8 79.9 77.6 70.8 79.9
-DuoAttention 92.8 91.9 89.0 85.0 81.1 76.1 86.0
-[…]
-Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
-Streaming 55.6 47.8 35.9 30.6 28.1 21.0 36.5
-Triangle 88.4 81.9 74.0 58.3 51.0 14.6 61.4
-StreamingMix 93.3 90.4 89.4 81.7 76.4 63.6 82.5
-[…]
-Table 3: Comparison with static sparsity methods on RULER.
-Llama-3.1, Llama-3-262K, Qwen2.5 are abbreviations for Llama-3.1-8B-Instruct, Llama-3-8B-262K, and Qwen2.5-8B-Instruct.
-The same applies to Table 4.
-Comparison with Static Sparsity.
-Table 3 and Table 2 present the evaluation results of TriangleMix and various static sparsity baselines.
-TriangleMix consistently outperforms all static baselines and remains nearly lossless performance across all input lengths.
-In contrast, other static methods usually suffer noticeable performance degradation.
-DuoAttention performs comparably to TriangleMix on LongBench using Llama-3.1-8B-Instruct.
-However, its performance is worse than TriangleMix on RULER, especially deteriorates at longer lengths (e.g., 64K and 128K).
-Furthermore, DuoAttention requires a separate training phase to learn head-wise sparsity patterns, which introduces additional computational overhead and may fail to converge on models such as Qwen-2.5-7B-Instruct.
-On the LongBench benchmark, we observe that the full Triangle pattern performs poorly on PassageRetrieval tasks, while the StreamingMix pattern underperforms on the in-context learning task (TREC).
-These findings suggest that attention over the Middle Q-K section in shallow layers is crucial for retrieval tasks, while attention over the Last Q-K section in deeper layers plays a key role in supporting in-context learning.
-Comparison with Dynamic Sparsity.
-Table 4 and Table 8 report results comparing TriangleMix with dynamic sparsity methods, as well as their combinations.
-FlexPrefill when $\gamma=0.90$ usually leads to performance degradation, and MInference generally outperforms it.
-While the performance of MInference and FlexPrefill with $\gamma=0.95$ is similar on the RULER benchmark, MInference still achieves better results on LongBench.
-When combined with either MInference or FlexPrefill, TriangleMix always fully retains the original performance of the dynamic sparsity methods.
-Moreover, combining TriangleMix with FlexPrefill even yields slight performance gains, particularly with Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct.
-Overall, TriangleMix preserves the full accuracy of dense attention across tasks while significantly reducing time complexity from $O(N^{2})$ to $O(N)$ for layers $l>L_{\text{tri\_start}}$.
-Methods 4K 8K 16K 32K 64K 128K Avg.
-Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
-MInference 96.3 95.1 95.0 90.5 86.8 75.0 89.8
-[…]
-FlexPrefill (0.90) 90.1 89.7 89.4 89.1 82.0 74.1 85.7
-[…]
-FlexPrefill (0.95) 95.0 94.7 94.5 92.8 86.5 75.9 89.9
-[…]
-Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
-MInference 93.4 90.7 88.8 85.1 82.6 80.1 86.8
-[…]
-FlexPrefill (0.90) 82.4 80.0 80.7 76.2 74.5 69.1 77.1
-[…]
-FlexPrefill (0.95) 90.3 87.7 88.0 83.8 80.2 75.7 84.3
-[…]
-Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
-MInference 95.6 93.8 92.8 84.9 79.1 68.0 85.7
-[…]
-FlexPrefill (0.90) 84.4 79.4 82.4 71.4 65.4 47.7 71.8
-[…]
-FlexPrefill (0.95) 91.1 89.7 88.5 75.4 72.5 51.2 78.1
-[…]
-Table 4: Comparison with dynamic sparsity methods on RULER.
-3.3 Analysis of Start Layer
-Figure 5: Average RULER score at 64K length for different $L_{\mathrm{tri\_start}}$ values.
-In this section, we investigate the choice of $L_{\text{tri\_start}}$ for each LLM.
-We evaluate all models on the 64K length tasks in the RULER benchmark, sweeping $L_{\text{tri\_start}}$ from 0 to the total number of layers with a step size of 2.
-As shown in Figure 5, Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit similar patterns: setting $L_{\text{tri\_start}}=12$ retains 99.7% and 100.0% of their original performance, respectively.
+3.3 Analysis of $L_{\mathrm{tri}}$
+Figure 4: Average RULER score at 64K length for different $L_{\mathrm{tri}}$ values.
+In this section, we investigate the choice of $L_{\text{tri}}$ for each LLM.
+We evaluate all models on the 64K length tasks in the RULER benchmark, sweeping $L_{\text{tri}}$ from 0 to the total number of layers with a step size of 2.
+As shown in Figure 4, Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit similar patterns: setting $L_{\text{tri}}=20$ retains 99.7% and 100.0% of their original performance, respectively.
 This indicates that 62.5% of the layers can adopt the $O(N)$ Triangle Attention without noticeable degradation.
-In contrast, Qwen-2.5-7B-Instruct requires a higher threshold: setting $L_{\text{tri\_start}}=20$ retains 97.9% of the performance, corresponding to a sparsity ratio of 28.6%.
-Method 32K 64K 128K
-Dense 44 179 750
-MInference 94 (0.5x) 152 (1.2x) 229 (3.3x)
-FlexPrefill (0.95) 32 (1.4x) 71  (2.5x) 220 (3.4x)
-FlexPrefill (0.90) 24 (1.8x) 52  (3.4x) 143 (5.2x)
-[…]
-Table 5: Attention kernel latency (ms) for one layer measured on Llama-3.1-8B-Instruct.
-Figure 6: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct.
+In contrast, Qwen-2.5-7B-Instruct requires a less threshold: setting $L_{\text{tri}}=8$ retains 98.2% of the performance, corresponding to a sparsity ratio of 28.6%.
+Llama-3.1-8B-Instruct Qwen2.5-7B-Instruct
+Method 32K 64K 128K 32K 64K 128K
+Dense 44 179 750 40 156 646
+MInference 94 (0.5x) 152 (1.2x) 229 (3.3x) 93 (0.4x) 151 (1.0x) 245 (2.6x)
+FlexPrefill 32 (1.4x) 71 (2.5x) 220 (3.4x) 38 (1.0x) 99 (1.6x) 310 (2.1x)
+XAttention 47 (0.9x) 127 (1.4x) 391 (1.9x) 44 (0.9x) 128 (1.2x) 391 (1.7x)
+Triangle 12 (3.7x) 24 (7.5x) 49 (15.3x) 11 (3.7x) 21 (7.4x) 76 (8.5x)
+Table 6: Attention kernel latency (ms) for one layer.
 Method 32K 48K 64K 80K 96K 112K 128K
 Dense 4.1 7.3 11.2 15.9 21.3 27.5 34.5
+DuoAttention 3.6 (-13%) 5.8 (-20%) 8.7 (-22%) 11.7 (-26%) 15.5 (-27%) 19.1 (-31%) 23.7 (-31%)
+TriangleMix 3.6 (-12%) 5.9 (-19%) 8.6 (-23%) 11.7 (-26%) 15.2 (-29%) 19.1 (-31%) 23.4 (-32%)
 MInference 5.5 (+34%) 7.8 (+7%) 10.1 (-10%) 12.3 (-23%) 13.4 (-37%) 15.9 (-42%) 18.0 (-48%)
-FlexPrefill (0.95) 3.6 (-12%) 5.5 (-25%) 7.7 (-31%) 9.9 (-38%) 12.3 (-42%) 15.0 (-45%) 17.8 (-48%)
-TriangleMix 3.6 (-12%) 5.9 (-19%) 8.6 (-23%) 11.7 (-26%) 15.2 (-29%) 19.1 (-31%) 23.4 (-32%)
 Ours + MInference 4.2 (+2%) 6.0 (-18%) 7.7 (-31%) 9.5 (-40%) 10.9 (-49%) 12.7 (-54%) 14.5 (-58%)
-Ours + FP (0.95) 3.4 (-17%) 5.2 (-29%) 7.2 (-36%) 9.2 (-42%) 11.3 (-47%) 13.6 (-51%) 16.1 (-53%)
-FlexPrefill (0.90) 3.3 (-20%) 5.1 (-30%) 6.8 (-39%) 8.9 (-44%) 11.1 (-48%) 13.6 (-51%) 15.5 (-55%)
-Ours + FP (0.90) 3.3 (-20%) 4.5 (-38%) 6.7 (-40%) 8.5 (-47%) 10.5 (-51%) 12.6 (-54%) 14.5 (-58%)
+FlexPrefill 3.6 (-12%) 5.5 (-25%) 7.7 (-31%) 9.9 (-38%) 12.3 (-42%) 15.0 (-45%) 17.8 (-48%)
+Ours + FlexPrefill 3.4 (-17%) 5.2 (-29%) 7.2 (-36%) 9.2 (-42%) 11.3 (-47%) 13.6 (-51%) 16.1 (-53%)
+XAttention 4.1 (-1%) 6.6 (-10%) 9.4 (-16%) 12.4 (-22%) 15.7 (-26%) 19.3 (-30%) 23.2 (-33%)
+Ours + XAttention 3.5 (-14%) 5.5 (-24%) 7.7 (-31%) 10.0 (-37%) 12.4 (-42%) 15.1 (-45%) 17.8 (-48%)
 Table 7: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct.
-We separate FlexPrefill ($\gamma=0.90$) as it generally leads to accuracy degradation, whereas the other settings do not.
 3.4 Efficiency of TriangleMix
-We implement Triangle Attention using Triton Tillet et al.
-(2019), with further implementation details provided in Appendix A.2.
+We implement Triangle Attention using Triton (Tillet et al., 2019), with further implementation details provided in Appendix A.2.
 All experiments are conducted on a single NVIDIA A100 80GB GPU.
-Dense attention is implemented using FlashAttention Kwon et al.
-(2023).
+Dense attention is implemented using FlashAttention (Kwon et al., 2023).
 Attention Kernel Latency.
-We first benchmark the average attention kernel time per layer on Llama-3.1-8B-Instruct at sequence lengths of 32K, 64K, and 128K.
-As shown in Table 5, Triangle Attention significantly outperforms dense attention, achieving 3.7×3.7$\times$3.7 × to 15.3×15.3$\times$15.3 × speedup.
+We first benchmark the average attention kernel time per layer on Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct at sequence lengths of 32K, 64K, and 128K.
+As shown in Table 6, Triangle Attention achieves $3.7\times$ to $15.3\times$ speedup compared to dense attention,.
 The speedup mainly comes from Triangle Attention’s linear complexity ($O(N)$), which makes it much more efficient than $O(N^{2})$ dense attention.
-Besides, unlike dynamic sparsity methods, it doesn’t require estimating attention block indices during runtime, which saves additional overhead.
+Besides, unlike dynamic sparsity methods, TriangleMix does not require estimating attention block indices.
+Its triangle attention kernel is also much simpler to implement, which further reduces overhead.
 End-to-End TTFT.
-We further measure time-to-first-token (TTFT) across sequence lengths of 32K, 48K, 64K, 96K, 112K, and 128K.
-Results for Llama-3.1-8B-Instruct are reported in Table 6.
-TriangleMix applies dense attention in shallow layers and Triangle Attention in deeper layers, reducing TTFT by 12%–32%.
-It is also compatible with dynamic attention for enhanced performance.
-For example, integrating MInference with TriangleMix reduces TTFT from 18.0s to 14.5s (a 19% decrease).
-Under the lossless setting, TriangleMix combined with FlexPrefill ($\gamma=0.95$) achieves the fastest TTFT for 32K–80K sequences, while TriangleMix with MInference is optimal for 96K–128K.
-We visualize the results of some selected methods in Figure 6.
-4 Discussion
-In this section, we discuss why Middle Q-K attention can be safely removed.
-Our hypothesis is that the observed phenomena arise from a Train–Test Misalignment.
-During training, the loss is applied uniformly to all positions in the input sequence.
-For example, if the input contains 2048 tokens, the model is also trained to predict token 1001 given tokens 1–1000.
-However, during inference, the primary goal is to predict tokens after the prompt (such as token 2049), making predictions of tokens inside the prompt (such as token 1001) largely irrelevant.
-To validate this assumption, we conduct an experiment on Llama-3.1-8B-Instruct.
-We sample 100 Wikipedia articles and extract the first 2048 tokens of each.
-In the last 16 layers, we apply Triangle Attention, skipping attention for Q indices between 512 and 1024.
-We then measure changes of the average perplexity (PPL) relative to dense attention in two output spans.
-Method PPL (1024–1152) PPL (1920–2048)
-Dense 8.02 7.79
-TriangleMix 8.34 (+0.32) 7.81 (+0.02)
-Table 7: With 2048 input tokens, we measure perplexity on two output spans.
-Triangle attention is applied, skipping attention between Q indices 512–1024 in the last 16 layers.
-Experiments are on Llama-3.1-8B-Instruct.
-The results in Table 7 show that skipping attention of Q index 512–1024 in deep layers noticeably increases perplexity when predicting tokens in the 1024–1152 range.
-For these positions, the skipped attention effectively becomes the Last Q-K attention necessary for predicting the next token.
-In contrast, for positions beyond 1920, the effect is minimal.
-The Middle Q-K region is important primarily for predicting tokens within the prompt, while for generation tasks focused on tokens after the prompt, the Middle Q-K in deeper layers is largely redundant.
-Although further theoretical work is needed, this observation suggests that selective sparsification of attention during inference can be done safely.
+We measure the end-to-end time-to-first-token (TTFT) for sequence lengths of 32K, 48K, 64K, 96K, 112K, and 128K.
+Table 7 reports the results on Llama-3.1-8B-Instruct.
+TriangleMix yields a TTFT reduction of 12%–32%.
+Compared with DuoAttention, TriangleMix achieves similar efficiency but causes less performance degradation (see Section 3.2).
+In addition, DuoAttention partitions heads within the same layer into different sparsity patterns (Liu et al., 2025), which leads to imbalance under tensor parallelism, while TriangleMix avoids this issue.
+Moreover, TriangleMix can be seamlessly combined with dynamic attention for further efficiency gains.
+For instance, integrating MInference with TriangleMix lowers TTFT from 18.0s to 14.5s (a 19% reduction) at length 128K.
+TriangleMix combined with FlexPrefill achieves the lowest TTFT for 32K–80K inputs, whereas TriangleMix with MInference is optimal for 96K–128K inputs.
 4 Related Works
 Static Sparsity Attention.
-Early methods employ fixed sparse patterns, such as strided Child et al.
-(2019), dilated Ding et al.
-(2023), sliding window Jiang et al.
-(2023), and mixed patterns Beltagy et al.
-(2020), typically requiring training from scratch.
-DuoAttention Xiao et al.
-(2024) introduces head-level static sparsity but needs additional offline training.
-Training-free approaches like Streaming Xiao et al.
-(2023) and Triangle Li et al.
-(2024a) attention often degrade accuracy for long contexts.
+Early methods employ fixed sparse patterns, such as strided (Child et al., 2019), dilated (Ding et al., 2023), sliding window (Jiang et al., 2023), and mixed patterns (Beltagy et al., 2020), typically requiring training from scratch.
+DuoAttention (Xiao et al., 2024) introduces head-level static sparsity but needs additional offline training.
+Streaming pattern is a training-free approach (Xiao et al., 2023) but leads to degraded accuracy for long contexts (Li et al., 2024a).
 In contrast, the proposed TriangleMix attention pattern significantly mitigates performance loss and is nearly lossless in accuracy.
 Dynamic Sparsity Attention.
-Methods like MInference Jiang et al.
-(2024a), FlexPrefill Lai et al.
-(2025), and XAttention Xu et al.
-(2025) dynamically determine attention masks online, maintaining high accuracy but incurring computational overhead for moderately long contexts.
-TriangleMix can enhance these dynamic approaches for further acceleration .
+Existing methods such as MInference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and XAttention (Xu et al., 2025) dynamically identify attention blocks with high scores and restrict computation to these blocks.
+While this form of attention score sparsity focuses on pruning blocks with small attention scores, our work moves beyond by uncovering a different kind of sparsity, which we term decoding-time contribution sparsity.
+Specifically, we show that many blocks with non-trivial scores still contribute little to decoding, and thus can be safely skipped.
+Leveraging this property enables additional acceleration in the prefilling stage.
 Long-context LLM Inference.
-FlashAttention Dao et al.
-(2022) speeds up attention by reducing memory access through fused operations.
-PagedAttention Kwon et al.
-(2023) improves decoding by managing KV cache allocation efficiently.
-KV cache optimizations techniques also include token-level eviction (SnapKV Li et al.
-(2024b)) and query-aware cache selection (Quest Tang et al.
-(2024)).
+FlashAttention (Dao et al., 2022) speeds up attention by reducing memory access through fused operations.
+PagedAttention (Kwon et al., 2023) improves decoding by managing KV cache allocation efficiently.
+KV cache optimizations techniques also include token-level eviction (SnapKV (Li et al., 2024b)) and query-aware cache selection (Quest (Tang et al., 2024)).
 TriangleMix is orthogonal to these approaches.
 5 Conclusion
-In this paper, we propose a hybrid attention pattern for long-context LLMs, where dense attention is used in shallow layers and Triangle attention is applied in deeper layers.
-This pattern is training-free and achieves nearly lossless accuracy.
-Our method achieves a 12% to 32% time-to-first-token reduction across 32K to 128K inputs.
-Moreover, it can be further accelerated when combined with dynamic sparsity attention.
-Limitations
-One limitation of our method is that the sparsity ratio is fixed within the pretrained model.
-We observe that Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit higher inherent sparsity compared to Qwen2.5-7B-Instruct, which constrains the achievable acceleration on the latter.
-It may be possible to enhance the sparsity of models like Qwen2.5-7B-Instruct through a post-training optimization process, which we leave for future exploration.
+In this paper, we identify a new form of sparsity in the prefilling stage, termed decoding-time contribution sparsity, and introduce TriangleMix, a training-free static attention pattern.
+By selectively applying Triangle attention in certain layers, TriangleMix substantially reduces attention overhead while maintaining nearly lossless performance.
+On 128K inputs, it achieves up to a 15.3× speedup in attention computation, surpassing typical dynamic sparse methods.
+Moreover, TriangleMix can be seamlessly combined with dynamic sparsity, yielding an additional 6%–19% reduction in TTFT.
 Appendix A Appendix
-A.1 Task to Probe Attention Importance
-We use a key-value retrieval task to analyze the importance of different attention sections.
-The language model is prompted to retrieve the correct value, beginning from its first generated token.
-The prompt is as follows:
+A.1 Details of Attention Block Probing
+We divide the causal attention matrix into three distinct sections as illustrated in Figure 1(a):
+  • Streaming section: includes attention sink and the sliding window;
+  • Last Q–K section: covers interactions between the last part of $\bm{Q}$ and $\bm{K}$, excluding the Streaming section;
+  • Middle Q–K section: consists of the remaining interactions between the middle parts of $\bm{Q}$ and $\bm{K}$.
+We define the Streaming mask as:
+$\bm{M}^{\mathrm{streaming}}_{i,j}=\begin{cases}1,&i\geq j,j\leq si\\ 1,&i\geq j,i-j\leq sl\\ 0,&\mathrm{otherwise}\end{cases}$
+where $si$ is the number of sink tokens and $sl$ is the sliding window size.
+We define the Last Q–K mask as:
+$\bm{M}^{\mathrm{last}}_{i,j}=\begin{cases}1,&i\geq j,N-i<last,\\ &j>si,i-j>sl\\ 0,&\mathrm{otherwise}\end{cases}$
+where $\textit{last}\geq 1$ specifies the number of rows corresponding to the last section.
+The Middle Q–K mask is defined as:
+$\bm{M}^{\mathrm{middle}}_{i,j}=\begin{cases}1,&i\geq j,N-i\geq last,\\ &j>si,i-j>sl\\ 0,&\mathrm{otherwise}\end{cases}$
+In our gradient-based attention block probing experiments, we set the parameters to $si=64$, $sl=128$, and $last=128$.
+We adopt a key-value retrieval task, where the language model is prompted to output the correct value starting from its very first generated token.
+Each input sequence contains approximately 2,000 tokens, and gradients are computed over 100 randomly generated samples.
+The prompt used in this task is as follows:
 […]
-We randomly generate pairs of keys and values using Version 4 UUID strings.
+We randomly generate pairs of keys and values using UUID strings.
 A.2 Implementation of Triangle Attention
-We implement Triangle Attention using Triton Tillet et al.
-(2019), with details provided in Algorithm 1.
+We implement Triangle Attention using Triton (Tillet et al., 2019), with details provided in Algorithm 1.
 For each row index, we apply a different attention mechanism: if the row index is less than $N-N_{\text{last}}$, we apply Streaming Attention; otherwise, we apply chunked FlashAttention to increase GPU utilization.
@@
  Calculate best number of splits $S$ for last row
- Initialize $\boldsymbol{O}\leftarrow(0)^{N\times d_{h}}$
- Initialize output buffer $\boldsymbol{O}_{\mathrm{last}}\leftarrow(0)^{SN_{\mathrm{last}}\times d_{h}}$
- Initialize LSE buffer $\boldsymbol{l}_{\mathrm{last}}\leftarrow(-\inf)^{SN_{\mathrm{last}}}$
+ Initialize $\boldsymbol{O}\leftarrow(0)^{N\times d_{h}}$, output buffer $\boldsymbol{O}_{\mathrm{last}}\leftarrow(0)^{SN_{\mathrm{last}}\times d_{h}}$, LSE buffer $\boldsymbol{l}_{\mathrm{last}}\leftarrow(-\inf)^{SN_{\mathrm{last}}}$
 Parallelized in GPU
 for $i\leftarrow 1$ to $\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil+S\lceil\frac{N_{\mathrm{last}}}{B_{M}}\rceil$ do
-  Initialize $\boldsymbol{O}_{\text{chip}}\leftarrow(0)^{B_{M}\times d_{h}}$
-  Initialize $\boldsymbol{m}\leftarrow(-\inf)^{B_{N}}$
-  Initialize $\boldsymbol{s}\leftarrow(0)^{B_{M}}$
+  Initialize $\boldsymbol{O}_{\text{chip}}\leftarrow(0)^{B_{M}\times d_{h}}$, $\boldsymbol{m}\leftarrow(-\inf)^{B_{N}}$, $\boldsymbol{s}\leftarrow(0)^{B_{M}}$
   if $i\leq\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil$ then
@@
    for $j\leftarrow 1$ to $\lceil\frac{N_{\mathrm{sink}}}{B_{N}}\rceil$ do
-    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$
-    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
+    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$, $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
     $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
@@
    for $j\leftarrow\lceil\frac{N-N_{\mathrm{window}}}{B_{N}}\rceil$ to $\lceil\frac{N}{B_{N}}\rceil$ do
-    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
-    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
+    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$, $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
     $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
@@
    for $j\leftarrow c\lceil\frac{N}{SB_{N}}\rceil$ to $(c+1)\lceil\frac{N}{SB_{N}}\rceil$ do
-    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$
-    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
+    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$, $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
     $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
@@
 $\boldsymbol{O}^{(N-N_{\mathrm{last}}):N}\leftarrow\mathrm{merge\_output}(\boldsymbol{O}_{\mathrm{last}},\boldsymbol{l}_{\mathrm{last}})$
-A.3 Additional Longbench Data
-Table 8 compares our method with dynamic sparsity attention on Longbench.
-When combined with MInference, our method achieves similar performance to MInference alone.
-Integrating with FlexPrefill yields a slight improvement over its original results.
-Methods Average
-2Wiki
-GovRep
-Hotpot
-LCC
-MNews
-MF-en
-PsgCnt
-PsgRtr
-Qasper
-Rbench
-Samsum
-TREC
-TrvQA
-Llama-3.1-8B-Instruct 54.8 48.1 34.4 61.6 66.6 25.6 56.9 16.8 99.7 44.9 52.7 42.6 71.0 91.6
-MInference 54.9 48.1 34.5 61.4 67.0 25.9 57.0 17.5 99.7 44.7 52.5 42.6 71.3 91.9
-[…]
-FlexPrefill; $\gamma=0.90$ 44.0 31.0 32.8 48.8 59.1 25.1 51.3 4.0 39.0 37.1 50.8 41.5 65.3 86.2
-[…]
-FlexPrefill; $\gamma=0.95$ 48.4 39.4 33.7 58.3 67.2 25.6 55.6 4.0 47.7 42.5 53.8 41.9 69.7 90.2
-[…]
-Llama-3-8B-Instruct-262k 44.2 21.0 34.3 26.4 46.1 26.3 50.2 4.7 95.0 30.5 43.3 41.1 68.3 86.8
-MInference 44.1 20.5 34.3 25.8 46.2 26.3 50.2 4.3 95.0 30.7 44.7 40.5 68.3 86.8
-[…]
-FlexPrefill; $\gamma=0.90$ 33.6 16.8 32.7 23.0 46.8 25.6 45.6 2.3 8.0 23.2 41.4 36.8 61.0 73.3
-[…]
-FlexPrefill; $\gamma=0.95$ 35.0 17.7 33.0 25.1 36.8 26.1 49.1 3.3 13.7 28.3 36.2 39.1 65.0 81.7
-[…]
-Qwen2.5-7B-Instruct 46.8 24.5 30.7 32.2 62.2 22.3 41.3 13.1 99.3 24.4 60.7 42.5 67.0 88.7
-MInference 46.8 24.5 30.7 32.7 61.6 22.3 40.8 12.8 100.0 24.3 60.2 42.1 68.0 89.3
-[…]
-FlexPrefill; $\gamma=0.90$ 37.3 17.5 29.3 27.6 64.0 21.8 39.0 7.0 28.7 15.1 51.4 39.3 64.7 79.3
-[…]
-FlexPrefill; $\gamma=0.95$ 40.8 20.2 29.7 33.9 66.8 21.8 40.4 4.0 48.7 17.1 54.8 40.5 67.3 85.5
-[…]
-Table 8: Comparison with dynamic sparsity methods on LongBench.
+A.3 Acknowledgement of LLM Usage
+We used large language models to polish the writing of this paper, and all generated content was carefully reviewed to ensure precise expression.